### v1

# custom.bib

Abstract
While Rotary Position Embeddings (RoPE) for natural language performs well and has become widely adopted, its adoption for other modalities has been slower. Here, we introduce Lie group Relative position Encodings (LieRE) that goes beyond RoPE in supporting higher dimensional inputs. We evaluate the performance of LieRE on 2D and 3D image classification tasks and observe that LieRE leads to marked improvements in performance (up to 6%), training efficiency (3.5x reduction), data efficiency (30%) compared to the baselines of RoFormer, DeiT III, RoPE-Mixed and Vision-Llama.
1 Introduction
While the attention mechanism has achieved widespread use, especially as part of the transformer architecture, attention is invariant to the order of its inputs and requires another mechanism to capture positional information of input tokens [vaswani2017attention]. This has spurred a line of work in the subarea of positional encodings—methods of encoding positional information for attention mechanisms.
In particular, Rotary Position Encoding (RoPE) has emerged as a powerful technique for encoding relative positional information in transformer-based models [RoFormer2024Su]. RoPE’s ability to capture relative position information has made it a popular choice for open-source foundation models such as LLaMA. In particular, RoPE implicitly captures relative positions. When the token in position $x$ attends to a token in position $y$, the effect of RoPE depends on $x-y$.
Despite the success of RoPE in sequence tasks, it is designed for one-dimensional sequence data. This has resulted in slow adoption for modalities with higher dimensional data, such as 2D images, and 3D medical images or data that includes a temporal dimension like videos or Electronic Health Records. Furthermore works that do attempt attempt to generalize RoPE to these modalities typically do so with a method tailored specifically to the modality at hand.
Our work aims to answer the question of whether there is a single relative position encoding scheme that well works across modalities of varying dimensions. If possible, this would enable the use of a common model backbone across a much larger variety of tasks.
$R(k_{1}x)$$R(k_{1}y)$$R(k_{2}x)$$R(k_{2}y)$
(a) 2D RoPE
$\exp(A\begin{bmatrix}x\\ y\end{bmatrix})$
(b) LieRE
Figure 1: Rotation matrices for a token generated from an image patch. $x,y$ denote the location of the patch from which the token was generated.
1.1 Contributions
We introduce Lie Relative Encodings (LieRE), a mechanism that allows the attention mechanism to learn how to utilize relative spatial information in its inputs. We show that LieRE is effective, without modification, across modalities and input dimensionality by evaluating it on both 2D and 3D classification tasks. Beyond improving classification accuracy, LieRE also reduces the amount of computing and data required during training. On the CIFAR100 task, this translates to 3.5x fewer training steps to achieve the same accuracy as the DeiT III baseline and outperforming the DeiT baseline trained on the full set with only a 70% subset of the data. Furthermore, LieRE is simple to implement and adapt to new modalities, requiring only a tokenizer that also outputs a position in $\mathbb{R}^{d}$ in addition to a standard embedding. In order to aid the reproducibility of our results we will post our code on github.
1.2 Related Work
1.2.1 Position Encodings
The development of various position encodings has been ongoing since the original development of the attention mechanisms. There are three broad classes of positional encodings: 1) absolute 2) relative and 3) contextual.
Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the text or media. Classic methods such as sinusoidal embeddings of learned embeddings achieve this by adding either a learned or carefully designed vector to the embedding of the token before it is passed through the transformer [vaswani2017attention, devlin2019bert, dosovitskiy2020image]. Absolute position measures position with respect to an absolute reference, such as the start of the text or the top left corner of an image.
Relative embeddings are focused on discarding the absolute coordinate system of absolute position encodings and encoding the relative positions of two tokens. This is particularly useful for images, where translation invariance can improve performance, but also widely used for text as well. The original relative position encodings modified the attention mechanism between two tokens with the difference in the positions of two tokens [shaw2018self, liu2021swin, liu2022swinv2]. However, this incurs quadratic computational complexity as an additional operation for every pair of tokens. Rotary Position Encodings (RoPE) avoid this penalty by utilizing the commutativity and orthogonality of 2D rotations. The keys and queries are rotated by a set of 2D rotations before computing the inner product, avoiding the quadratic blowup of earlier relative position schemes. The angle of these rotations depends on the position of the original token. The algebraic properties of attention and 2D rotation matrices ensure that the attention mechanism only depends on the relative positions of the tokens when they attend to each other [RoFormer2024Su]. RoPE is quite widely used in open source LLMs including the PaLM, Llama and Mixtral models [touvron2023llama, chowdhery2023palm, jiang2024mixtral]. However, RoPE can perform poorly on inference for larger context sizes than the model was trained on.
We refer to the last category of positional as contextual position embeddings. This category is defined by encodings that aim to capture semantic positional information lost in traditional absolute and relative position encodings, often motivated by reasoning or mathematical tasks. Contextual Position encodings achieve (CoPE) this by allowing the model to learn how the position is computed [golovneva2024contextual]. Abacus embeddings enable transformers to learn how to handle arithmetic by better exposing the digit structure of numbers [mcleish2024transformers].
1.2.2 Extensions of RoPE
The efficiency and popularity of RoPE have led to several lines of work looking to extend it focused on specific domains.
One notable one is context extension, which aims to address the fact that models trained on short documents tend to perform poorly on long documents. Methods like NTK-aware context extension, YaRN and LongROPE focus on enabling already trained models to handle long context, both with and without finetuning [ding2024longrope, peng2023yarn].
Another line of work has been specifically focused on adapting RoPE to image tasks. Both VisionLlama and RoPE-Mixed present relative position encodings inspired by RoPE that are able to encode 2D positional data [chu2024visionllama, heo2024rotary] I key difference is that RoPE-Mixed has a learnable component, whereas VisionLlama does not.
2 Background
2.1 Lie Groups in the Context of Attention
In this section, we aim to provide a minimal introduction to Lie groups so that the reader is able to understand the mathematical motivations behind LieRE. Lie groups are very well studied, especially in the context of representation theory and standard texts including [fulton2013representation] are able to provide a more extensive introduction to the subject.
In this context, Lie groups are smooth manifolds that are closed under matrix multiplication and inversion. For every Lie group, the matrix exponential provides a smooth bijective map from a subset of $\mathbb{R}^{n\times n}$ (hereto referred to as the generator set) to the Lie group. The exponential map is a diffeomorphism and has the following key property for generators $x,y$ close together:
$\exp(x-y)=\exp(-y+x)\approx\left(\exp(y)\right)^{-1}\exp(x)$ (1)
Section 3.1 will show that this is the key to being able to encode relative positions.
Both RoPE (in the context of text) and RoPE-Mixed use block-diagonal rotation matrices with 2D rotations as blocks. These form a special Lie group that is commutative, allowing us to strengthen the statement in equation (1) to
$exp(x-y)=exp(x)exp(y)^{-1}=exp(y)^{-1}exp(x).$ (2)
Our work examines the tradeoff between using the stronger property in equation (2) or increased capacity and the weaker property (1). In short, we find that benefit of representational flexibility exceeds the benefit of exactly honoring equation (2).
3 Methods
3.1 Rotations for Relative Positions in Attention
Before introducing the pseudocode we explain how LieRE modifies the attention mechanism and how this relates to Lie groups.
We start by using a linear map to embed the positions as skew-symmetric matrices $P_{i}$ and $P_{j}$ and compute high dimensional rotations as $R_{i}=\exp(P_{i})$ and $R_{j}=\exp(P_{j})$.
Recall that, for every pair of tokens, the attention mechanism computes the inner products between their key and query vectors, $k_{i}^{T}q_{j}$. We encode the positions by multiplying by the rotation matrices in the prior part. In particular, we update the keys and queries as $k_{i}^{\prime}=k_{i}R_{i}$ and $q_{j}=q_{j}R_{j}$. This results in an updated inner product of $(R_{i}k_{i})^{T}R_{j}q_{j}=k_{i}R_{i}^{T}R_{j}q_{j}=k_{i}R_{i}^{-1}R_{j}q_{j}.$
Recall that by equation 1, $R_{i}^{-1}R_{j}=\exp(P_{i})^{-1}\exp(P_{j})\approx\exp(P_{j}-P_{i})$. In other words, the inner product automatically computes the relative position encoding. Note that this is exactly the mechanism behind the original RoPE paper. The difference is that LieRE uses the full set of high dimensional rotations and relaxes the constraint that equation 1 holds with equality. We make the description of the method more precise in the next section.
3.2 LieRE
LieRE modifies the attention mechanism by applying a rotation to the keys and queries before computing their inner products. However, where RoPE applies a block-diagonal $n\times n$ with 2D rotation matrices as blocks, LieRE applies a $n\times n$ rotation matrix. Using the fact that every rotation matrix can be represented as the matrix exponential of a skew-symmetric matrix, we parametrize the rotations with generators before the matrix exponential. This skew-symmetric generator is computed via a learned linear map from the positions to the generator space. Let $x\in\mathbb{R}^{n}$ indicate the position of the token, and $A:\mathbb{R}^{d}\rightarrow\text{Skew}_{n}(\mathbb{R})$ be a learnable linear map to the set of skew-symmetric matrices. At the start of every forward pass, we compute the LieRE rotation matrix for each token as
$R_{LieRE}:=\exp(Ax).$
Then, for each attention mechanism, the keys and queries for a particular token are updated as $\mathrm{key^{\prime}}:=R_{LieRE}\cdot\mathrm{key},\mathrm{query^{\prime}}:=R_{LieRE}\cdot\mathrm{query}$ We present attention with RoPE and LieRE side by side in algorithms 1 and 2.
Algorithm 1 RoPE Attention (prior work)
1:procedure RoPE($X,p,d$)
2:     $\theta\leftarrow\frac{1}{10000^{2i/d}}\quad\forall i\in[0,d)$
3:     for $i\leftarrow 0$ to $d$ step $2$ do
[…]
5:         $X_{\text{rotated}}[i]\leftarrow X[i]\cdot\text{cos\_pos}_{i}-X[i+1]\cdot\text{sin\_pos}_{i}$
6:         $X_{\text{rotated}}[i+1]\leftarrow X[i]\cdot\text{sin\_pos}_{i}+X[i+1]\cdot\text{cos\_pos}_{i}$
7:     end for
8:     return $X_{\text{rotated}}$
9:end procedure
10:procedure RoPEAttention($Q,K,V$)
11:     $p\leftarrow\text{tokenPositions}$
[…]
13:     $K_{\text{rotated}}\leftarrow\textsc{RoPE}(K,p,d)$
14:     $Q_{\text{rotated}}\leftarrow\textsc{RoPE}(Q,p,d)$
15:     $\text{Attention}\leftarrow\text{softmax}(\frac{Q_{\text{rotated}}K_{\text{rotated}}^{T}}{\sqrt{d}})V$
16:     return Attention
17:end procedure
Algorithm 2 LieRE Attention (our work)
1:procedure LieRE($X,p,A$)
[…]
4:     return $X_{\text{rotated}}$
5:end procedure
6:procedure LieREAttention($Q,K,V,A$)
7:     $p\leftarrow\text{tokenPositions}$
8:     $K_{\text{rotated}}\leftarrow\textsc{LieRE}(K,p,A)$
9:     $Q_{\text{rotated}}\leftarrow\textsc{LieRE}(Q,p,A)$
10:     $\text{Attention}\leftarrow\text{softmax}\left(\frac{Q_{\text{rotated}}K_{\text{rotated}}^{T}}{\sqrt{\text{dim}(K)}}\right)V$
11:     return Attention
12:end procedure
Outside of the generator scaling experiments, we set the rotation dimension equal to the head dimension and apply rotation on a per-attention-head basis.
Note that if we are applying our method to images it is possible to impose a sparsity structure on the generator that recovers RoPE-mixed without the absolute position embeddings [heo2024rotary].
4 Experiments
4.1 Model and Training Parameters
In order to isolate the benefit of LieRE, we use the standard transformer backbone modified to be able to switch between relative position encoding types. The backbone is configured as ViT-B, with 12 layers, a hidden dimension of 768, and an intermediate dimension of 3096. All experiments use dropout of 0.1. All experiments use RandAugment [cubuk2020randaugment].
We avoid using pre-trained weights in order to maximize the comparability of results between methods. All trainers use ADAM with the default parameters for pytorch.
In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters with LieRE and use the same default hyperparameters for all experiments (Appendix A).
4.2 Datasets and Tasks
Our experiments are designed to evaluate the efficacy of LieRE as a position encoding across a diverse set of modalities. We evaluate LieRE on the classification of 2D (images) and 3D (videos and medical images) data. For 3D data and ImageNet (2D), we focus on accuracy. For CIFAR100 (2D), where training is less resource intensive, we also evaluate LieRE’s data and training compute efficiency.
4.2.1 2D Classification
For 2D data we evaluate performance on the CIFAR100 and ImageNet image classification task [krizhevsky2009learning, deng2009imagenet]. All hyperparameters and other details necessary for the reproducibility of experiments can be found in appendix A.
We partition our evaluation of performance into four parts. In the first part, we examine accuracy across a range of model architectures on both CIFAR100 and ImageNet. To ensure a fair comparison, we implement all positional encodings on the same [standard] transformer backbone. These experiments are meant to identify the best-performing method of encoding positions.
In the second part, we take advantage of the relatively modest amount of compute resources necessary to train a model for CIFAR 100 to examine LieREs impact on data efficiency. Here we limit our focus to LieRE, RoPE-Mixed, VisionLlama embeddings and absolute position embeddings. We train every model on 30%, 40%,…,90% subsets of CIFAR100 in order to evaluate the sensitivity of each model to the amount of training data. Since training starts to overfit quite early for the smaller datasets, we use the maximum validation accuracy for each run when comparing models.
We also measure training compute efficiency by comparing the number of training steps necessary to achieve a fixed level of validation accuracy.
In the third part, we evaluate the scaling laws of the LieRE generator capacity by imposing a block diagonal structure on the generator that allows us to vary the added capacity. With a head dimension of $64$ LieRE adds $4032$ parameters to a transformer designed for $2D$ data. While this is a small number of parameters, they lie in a critical point in the architecture, and understanding the marginal impact of this capacity can help inform computational tradeoffs.
The block diagonal generator structure results in a block-diagonal rotation matrix after the matrix exponential. The parameters for each block are learned separately. This setup allows us to vary the capacity of LieRE without modifying the rest of the model.
In the last part, we measure how much different models depend on the positional information in the image by shuffling the patches. A higher accuracy drop means the model relies more on the positions of the patches during inference.
4.2.2 3D Classification
In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-mixed and absolute encoding similar to the previous section [arnab2021vivit, heo2024rotary].
For the 3D experiments we examine video classification performance in the UCF101 dataset [soomro2012ucf101] and hemorrhage classification on the RSNA dataset [rsna2019data]. We treat the RSNA dataset as a binary classification problem, asking the model to identify whether a 3D brain scan has a hemorrhage or not.
Again, we did not optimize any hyperparameters for the LieRE model. We additionally use Mixup described standard used augmentation method for videos [zhang2017mixup]. The full set of hyperparameters may be found in appendix A.
5 Results
5.1 Accuracy
For 2D images, we show that the LieRE-based transformer outperforms the DeiT by 5.5% [touvron2022deit], RoPE adaptation in VisionLlama [chu2024visionllama] by 3.6% and RoPE-Mixed by 2.7% on Cifar100 [heo2024rotary].
For 3D inputs (video and 3D medical images), we observe an accuracy improvement of LieRE-based transformer over the Absolute Position Embedding of up to 6.7 % and RoPE-based transformers by up to 2.5% in accuracy.
Table 1: 2D image classification on Cifar100 and ImageNet [krizhevsky2009learning, deng2009imagenet], equivalent to DeiT
Method # params Cifar100 Accuracy (%) ImageNet Accuracy (%)
Absolute Position E. 85.1M 63.9 66.1
VisionLlama RoPE 85.1M 65.8 -
RoPE-Mixed 85.1M 66.7 68.5
LieRE (ours) 85.1M 69.4 68.8
Table 2: 3D video classification on RSNA hemorrhage prediction and UCF101 dataset [soomro2012ucf101, rsna2019data], equivalent to Vivit (spatio-temporal)
Method # params UCF101 Accuracy (%) RSNA Accuracy (%)
Absolute Position E. 88.7M 44.4 80.7
RoPE-Mixed 88.7M 48.6 81.9
LieRE (ours) 88.7M 51.1 82.7
5.2 Data efficiency
We observe that transformers based on LieRE exhibit greater data efficiency compared to leading transformer architectures for 2D images on the CIFAR-100 dataset. For instance, using only 40% of the dataset, the LieRE-based transformer outperforms all comparable methods by 4.47%. Additionally, there is a mild trend indicating that the performance gap widens as the amount of data decreases (Figure 2(b)).
(a) Performance scales with generator dimension (learnable parameters)
(b) Data ablation on Cifar100
Figure 2:
5.3 Generator Dimension Scaling Laws
LieRE adds a small amount of capacity to the model ($4032$ parameters in our case) but still represents the largest amount of learnable capacity around relative position encodings to date. This leads to the natural question of how helpful marginal capacity is.
We vary capacity by enforcing a block diagonal structure on the generator. varying the block size allows us to approach "full" LieRE. Note that $2\times 2$ blocks correspond exactly to RoPE-Mixed.
In figure 2(a) we evaluate accuracy versus block dimension. We observe the steepest increase at the very start, though accuracy continues to grow as the generator dimension increases.
5.4 Compute efficiency
Figure 3: The LieRE spatial encoding allows the model to match the performance of absolute position encodings with substantially less training time.
Training transformers can necessitate substantial computational resources, which can hinder equitable access to research and development of machine learning methods. We demonstrate that the LieRE-based transformer requires 3.5 times less training time to achieve comparable performance to the Absolute Position Embedding baseline (as used in DeiT III [touvron2022deit]). This represents the largest reduction in training time compared to recent works such as VisionLlama and RoPE-Mixed [chu2024visionllama, heo2024rotary]. We note that this is the first training efficiency comparison of these recent methods. Figure 3 shows the amount of allowable compute reduction to achieve the same accuracy achieved by absolute position encodings (DEIT baseline) after 200 epochs. LieRE demonstrates the largest win, allowing a 3.5X reduction in training compute to achieve the same accuracy.
While all methods display better performance with more training compute, a $3.5x$ reduction allows for significantly faster experimentation and makes a fixed level of performance more accessible to those with limited access to compute resources.
5.5 Patch shuffling
Shuffling patches and frames allows us to see how much the model is able to use the positional information in its inputs. A model whose architecture does not allow/encourage the use of positional information will converge to a representation similar in spirit to a bag-of-words, where the relative locations of pixels/voxels do not matter. A greater dropoff in accuracy during shuffling is indicative that the model more heavily utilizes positional information.
We evaluate models using the decline in accuracy when evaluating on shuffled patches. We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE encodes meaningful relative positions between tokens. The complete results are displayed in table LABEL:tab:shuffling_results for cifar100 and table LABEL:tab:shuffling_results_3D for UCF101.
Table 3: Relative drop in accuracy for 2D image classification after shuffling patch order, Cifar100
Method before shuffling $\uparrow$ after shuffling $\downarrow$ Accuracy Drop (%) $\uparrow$
Absolute Position E. 63.9 19.6 69.3
VisionLlama RoPE 65.8 29.7 54.8
RoPE-Mixed 66.7 16.5 75.4
LieRE (ours) 69.4 15.9 77.1
Table 4: Accuracy for Video recognition after patch shuffling, UCF101 dataset
Method before shuffling $\uparrow$ after shuffling $\downarrow$ Accuracy Drop (%) $\uparrow$
Absolute Position E. 44.4 44.0 0.1
RoPE-Mixed 48.6 36.6 24.7
LieRE (ours) 51.1 32.2 36.9
6 Limitations
While LieRE shows promising results across multiple modalities and input dimensionalities, there are a few limitations worth noting:
1. Compatibility with architectures that do not use attention: Our method is specifically designed to modify the inner product, making it compatible with most attention schemes, including original attention and linear attention. However, this specificity may limit its applicability to other architectures, like convolutional neural networks, that do not rely on inner product-based attention mechanisms. Future work could explore adaptations of our method to a wider range of architectures.
2. Limited to positions in $\mathbb{R}^{n}$: The current formulation of our method is designed to encode positions in $\mathbb{R}^{n}$. While this is sufficient for many applications, it may not be directly applicable to tasks that require encoding poses in $SE(3)$, like robotics. Further research may be necessary to adapt our method to effectively handle such representational requirements.
Despite these limitations, we believe that our work provides much-needed insight into how to improve model performance and reduce training costs by encoding relative position information across various modalities and input dimensionalities.
7 Broader Impacts
LieRE, our proposed method for encoding positional information in attention mechanisms, has demonstrated substantial improvements in 2D and 3D image classification tasks. We are particularly excited about how LieRE can expand the applicability of RoPE-based transformers to $n$ dimensions. This lays down the tracks to apply that same relative position within and across modalities and settings.
1. One Backbone for All Modalities: LieRE enhances the generalizability of attention mechanisms, making them more versatile and adaptable to a wider range of modalities. Once the input is encoded as tokens and positions, no additional changes are necessary in order to support a new modality. By providing a flexible and efficient way to encode positional information, LieRE can facilitate the application of attention mechanisms to diverse data types, such as audio, video, multi-dimensional sensor data or temporal data with diverse time intervals. This increased generalizability can lead to more effective transfer learning and cross-pollination of ideas between different research areas, accelerating progress and innovation in the field.
2. Improving accessibility of attention-based architectures for low-data and low-compute regimes: This is particularly valuable for applications where labeled spatial data is scarce or expensive to acquire, such as in medical imaging and robotics. By lowering the barriers to entry, LieRE can enable the application of attention mechanisms to a broader set of problems, including those in resource-constrained settings or niche domains.
3. Enabling more efficient and sustainable AI: The reduced training costs and improved data efficiency associated with LieRE can contribute to the development of more efficient and sustainable AI systems. By requiring fewer computational resources and less training data, LieRE can help mitigate the environmental impact of large-scale model training. This is particularly important as the field of AI continues to grow and the demand for computational resources increases.
In conclusion, LieRE has the potential to expand the generalizability of attention mechanisms, improve their accessibility for spatial data, and contribute to the development of more efficient and sustainable AI systems.
8 Conclusion
In this paper, we introduced Lie group Relative position Encodings (LieRE), a position encoding that can effectively encode relative position information for attention mechanisms across modalities and input dimensionalities. Through extensive experiments on 2D image classification (CIFAR100, ImageNet) and 3D medical image classification (RSNA brain CT hemorrhage detection), we demonstrated that LieRE achieves superior performance compared to existing positional encoding methods. Beyond improving accuracy, LieRE also exhibits data and compute efficiency. On CIFAR100, LieRE requires 3.5 times less training computed to match the performance of the baseline model with absolute position encodings. Furthermore, LieRE can outperform the baseline trained on the full dataset while using only 70% of the training data, highlighting its data efficiency. The key advantages of LieRE include its simplicity, flexibility, and ease of adaptation to new modalities. By requiring no changes to the tokenizer other than outputting positions in $\mathbb{R}^{n}$ and no other code changes, LieRE provides a unified and efficient approach for transformers to process and learn from various data modalities within a single architecture.
We hope that this empowers the field in building models that can understand concepts.
Funding and Acknowledgments We would like to thank Maksim Maydanskiy for helping us understand Lie groups and Lie group representations. We would like to thank Aradhana Sinha for suggesting the shuffling experiment and providing early feedback on the paper.
Sophie Ostmeier was supported by the German Research Foundation (DFG), Walter-Benjamin fellowship (ID: 517316550).
This project was supported by Google Cloud and Azure credits.
[…]
Appendix A Experiment Hyperparameters
The backbone for all experiments is configured as ViT-B, with 12 layers, a hidden dimension of 768, and an intermediate dimension of 3096. We use dropout of 0.1.
A.1 2D Image Classification
The CIFAR experiments where trained on 8xL4 GPUs with 24GB of VRAM each and all took under 30 minutes to complete. The generator scaling experiment was conducted using RTX6000 GPUs. The ImageNet experiments were trained on 8xL40 GPUs and all took less than 2 days and 5 hours of runtime including time lost due to preemption and resource sharing. We use a cosine learning rate schedule with an initial learning rate of $41E-$ 4 and train for 200 epochs. We use an effective batch size of $512$. We use a patch size of $4\times 4$ on the original $32\times 32$ image for Cifar100 and a patch size of $16\times 16$ on the randomly cropped and resized $224\times 224$ image. All vision experiments used RandAugment [cubuk2020randaugment]. We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$. The hyperparameters were tuned with RoPE-mixed and selected before conducting the LieRE trainers as to ensure a fair comparison.
A.2 3D Medical Image and Video Classifications
The 3D classification experiments were conducted on either $1008\times A$ 100 40GB GPUs or $1004\times A$ 100 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size. Similar to 2D classification, we use an initial learning rate of $41E-$ 4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $4\times 16\times 16$ on the randomly cropped and resized $32\times 224\times 224$ video/image. We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$. We use Mixup for all 3D experiments [tong2022videomae, zhang2017mixup].

### v2

@@
 Abstract
-While Rotary Position Embeddings (RoPE) for natural language performs well and has become widely adopted, its adoption for other modalities has been slower.
-Here, we introduce Lie group Relative position Encodings (LieRE) that goes beyond RoPE in supporting higher dimensional inputs.
-We evaluate the performance of LieRE on 2D and 3D image classification tasks and observe that LieRE leads to marked improvements in performance (up to 6%), training efficiency (3.5x reduction), data efficiency (30%) compared to the baselines of RoFormer, DeiT III, RoPE-Mixed and Vision-Llama.
+While Rotary Position Embeddings (RoPE) for large language models have become widely adopted, their application for other modalities has been slower.
+Here, we introduce Lie group Relative position Encodings (LieRE) that goes beyond RoPE in supporting n-dimensional inputs.
+We evaluate the performance of LieRE on 2D and 3D image classification tasks and observe that LieRE leads to marked relative improvements in performance (up to 9.7% for 2D and up to 25.5% for 3D), training efficiency (3.5x reduction), data efficiency (30%) compared to the baselines of DeiT III, RoPE-Mixed and Vision-Llama.
+https://github.com/Stanford-AIMI/LieRE
 1 Introduction
 While the attention mechanism has achieved widespread use, especially as part of the transformer architecture, attention is invariant to the order of its inputs and requires another mechanism to capture positional information of input tokens [vaswani2017attention].
-This has spurred a line of work in the subarea of positional encodings—methods of encoding positional information for attention mechanisms.
-In particular, Rotary Position Encoding (RoPE) has emerged as a powerful technique for encoding relative positional information in transformer-based models [RoFormer2024Su].
-RoPE’s ability to capture relative position information has made it a popular choice for open-source foundation models such as LLaMA.
+This has spurred a line of work in the subarea of positional encodings—methods of encoding positional information in attention mechanisms.
+In particular, Rotary Position Encoding (RoPE) has emerged as a technique for encoding relative positional information of text tokens in transformer-based models [RoFormer2024Su].
+RoPE’s ability to capture relative position information has made it a popular choice for open-source language foundation models such as LLaMA.
 In particular, RoPE implicitly captures relative positions.
-When the token in position $x$ attends to a token in position $y$, the effect of RoPE depends on $x-y$.
-Despite the success of RoPE in sequence tasks, it is designed for one-dimensional sequence data.
-This has resulted in slow adoption for modalities with higher dimensional data, such as 2D images, and 3D medical images or data that includes a temporal dimension like videos or Electronic Health Records.
-Furthermore works that do attempt attempt to generalize RoPE to these modalities typically do so with a method tailored specifically to the modality at hand.
-Our work aims to answer the question of whether there is a single relative position encoding scheme that well works across modalities of varying dimensions.
-If possible, this would enable the use of a common model backbone across a much larger variety of tasks.
-$R(k_{1}x)$$R(k_{1}y)$$R(k_{2}x)$$R(k_{2}y)$
-(a) 2D RoPE
-$\exp(A\begin{bmatrix}x\\ y\end{bmatrix})$
-(b) LieRE
-Figure 1: Rotation matrices for a token generated from an image patch. $x,y$ denote the location of the patch from which the token was generated.
+When the token in position $p_{1}$ attends to a token in position $p_{2}$, the effect of RoPE depends on $p_{1}-p_{2}$.
+Despite the success of RoPE in sequence tasks [touvron2023llama, chowdhery2023palm], it is designed for one-dimensional sequence data.
+This has resulted in slow adoption for modalities with higher dimensional data, such or data that includes a temporal dimension like videos.
+Our work aims to answer whether a single position encoding scheme can work well across both 2D and 3D modalities.
+If possible, this would enable the use of a simpler common model backbone across a variety of tasks.
+[…]
+Figure 1: LieRE sketch, where $A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x\;y\;z])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the first patch in the top left corner of the input with the position $P_{i}=(x,y,z)\in\mathbb{R}^{N}$.
+$d$ is the head dimension.
+The upper triangle contains learnable $\textrm{LieRE}_{\theta}$ parameter, while the lower triangle contains their negatives, -$\textrm{LieRE}_{\theta}$, reflecting the skew-symmetric nature of the matrix.
+[…]
+Figure 2: Comparisons of basis matrix $A$ of our work, Rope-Mixed [heo2024rotary], VisionLlama [chu2024visionllama].
+First row: learnable parameter (yellow), not learnable parameter (blue) in basis matrix.
+Second row: rotation matrices $\exp(Ax)$ shared across the stem (gray) or different for each layer and head (colorful).
+MHA := Multi-Head-Attention, FF := Feed Forward, L[index] := Layer index, RxH := Rotation Matrix per Head, $\textrm{LieRE}_{\theta}$ := the number of LieRE parameters.
 1.1 Contributions
-We introduce Lie Relative Encodings (LieRE), a mechanism that allows the attention mechanism to learn how to utilize relative spatial information in its inputs.
-We show that LieRE is effective, without modification, across modalities and input dimensionality by evaluating it on both 2D and 3D classification tasks.
-Beyond improving classification accuracy, LieRE also reduces the amount of computing and data required during training.
-On the CIFAR100 task, this translates to 3.5x fewer training steps to achieve the same accuracy as the DeiT III baseline and outperforming the DeiT baseline trained on the full set with only a 70% subset of the data.
-Furthermore, LieRE is simple to implement and adapt to new modalities, requiring only a tokenizer that also outputs a position in $\mathbb{R}^{d}$ in addition to a standard embedding.
+We introduce Lie Relative Encodings (LieRE), a mechanism that allows the attention mechanism to learn how to utilize relative spatial information of its inputs.
+We show that LieRE is effective on both 2D and 3D inputs of various modalities.
+Beyond improving classification accuracy, LieRE also reduces the amount of compute and data required during training to achieve a fixed accuracy.
+On the CIFAR100 task, this translates to 3.5x fewer training steps to achieve the same accuracy as the DeiT baseline and outperforming the DeiT baseline trained on the full set with only a 70% subset of the data.
+Furthermore, LieRE is simple to implement and adaptable to modalities, requiring only a tokenizer that also outputs a position in $\mathbb{R}^{d}$ in addition to a standard embedding.
 In order to aid the reproducibility of our results we will post our code on github.
@@
 1.2.1 Position Encodings
-The development of various position encodings has been ongoing since the original development of the attention mechanisms.
-There are three broad classes of positional encodings: 1) absolute 2) relative and 3) contextual.
-Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the text or media.
-Classic methods such as sinusoidal embeddings of learned embeddings achieve this by adding either a learned or carefully designed vector to the embedding of the token before it is passed through the transformer [vaswani2017attention, devlin2019bert, dosovitskiy2020image].
+The fact that the original attention mechanism is invariant to the order of tokens has motivated the ongoing development of methods to incorporate positional information into the transformer architecture.
+We split our literature review into three broad classes of positional encodings: 1) absolute, 2) relative, and 3) contextual.
+Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the intput.
+Methods such as sinusoidal and learned absolute encodings add vectors to the input token embedding [vaswani2017attention, devlin2019bert, dosovitskiy2020image].
 Absolute position measures position with respect to an absolute reference, such as the start of the text or the top left corner of an image.
-Relative embeddings are focused on discarding the absolute coordinate system of absolute position encodings and encoding the relative positions of two tokens.
-This is particularly useful for images, where translation invariance can improve performance, but also widely used for text as well.
-The original relative position encodings modified the attention mechanism between two tokens with the difference in the positions of two tokens [shaw2018self, liu2021swin, liu2022swinv2].
-However, this incurs quadratic computational complexity as an additional operation for every pair of tokens.
-Rotary Position Encodings (RoPE) avoid this penalty by utilizing the commutativity and orthogonality of 2D rotations.
-The keys and queries are rotated by a set of 2D rotations before computing the inner product, avoiding the quadratic blowup of earlier relative position schemes.
-The angle of these rotations depends on the position of the original token.
-The algebraic properties of attention and 2D rotation matrices ensure that the attention mechanism only depends on the relative positions of the tokens when they attend to each other [RoFormer2024Su].
+Relative position encodings instead encode the relative positions of two tokens.
+One strategy is to learn an embedding for position deltas which can be incorporated into the attention mechanism [shaw2018self, liu2021swin, liu2022swinv2].
+However, this incurs quadratic computational cost in terms of the number of tokens.
+Rotary Position Encodings (RoPE) avoid this cost by rotating the key and query vectors before the attention inner product.
+The algebraic properties of the block-diagonal rotation matrices used in RoPE ensures only relative positional information is captured in the attention mechanism [RoFormer2024Su].
 RoPE is quite widely used in open source LLMs including the PaLM, Llama and Mixtral models [touvron2023llama, chowdhery2023palm, jiang2024mixtral].
 However, RoPE can perform poorly on inference for larger context sizes than the model was trained on.
+This has spurred an active line of work extending RoPE to longer contexts, work which we review later.
 We refer to the last category of positional as contextual position embeddings.
@@
 The efficiency and popularity of RoPE have led to several lines of work looking to extend it focused on specific domains.
-One notable one is context extension, which aims to address the fact that models trained on short documents tend to perform poorly on long documents.
-Methods like NTK-aware context extension, YaRN and LongROPE focus on enabling already trained models to handle long context, both with and without finetuning [ding2024longrope, peng2023yarn].
+One notable one is context extension, which aims to address the fact that RoPE NLP models trained on short documents tend to perform poorly on long documents.
+Methods like NTK-aware context extension, YaRN and LongRoPE focus on enabling already trained models to handle long context, both with and without finetuning [ding2024longrope, peng2023yarn, tworkowski2024focused, chen2023extending].
 Another line of work has been specifically focused on adapting RoPE to image tasks.
-Both VisionLlama and RoPE-Mixed present relative position encodings inspired by RoPE that are able to encode 2D positional data [chu2024visionllama, heo2024rotary] I key difference is that RoPE-Mixed has a learnable component, whereas VisionLlama does not.
+Both VisionLlama and RoPE-Mixed present relative position encodings inspired by RoPE that are able to encode 2D positional data [chu2024visionllama, heo2024rotary].
+The primary difference is that RoPE-Mixed has a learnable component, whereas VisionLlama does not.
 2 Background
@@
 In this section, we aim to provide a minimal introduction to Lie groups so that the reader is able to understand the mathematical motivations behind LieRE.
-Lie groups are very well studied, especially in the context of representation theory and standard texts including [fulton2013representation] are able to provide a more extensive introduction to the subject.
-In this context, Lie groups are smooth manifolds that are closed under matrix multiplication and inversion.
+Lie groups are well studied, especially in the context of representation theory, and standard texts including [fulton2013representation] are able to provide a more extensive introduction to the subject.
+In this context, Lie groups are smooth sets of matrices that are closed under matrix multiplication and inversion.
 For every Lie group, the matrix exponential provides a smooth bijective map from a subset of $\mathbb{R}^{n\times n}$ (hereto referred to as the generator set) to the Lie group.
-The exponential map is a diffeomorphism and has the following key property for generators $x,y$ close together:
-$\exp(x-y)=\exp(-y+x)\approx\left(\exp(y)\right)^{-1}\exp(x)$ (1)
+The exponential map is a diffeomorphism and has the following key property for generators $U,V\in\mathbb{R}^{n\times n}$ close together:
+$\exp(U-V)=\exp(-V+U)\approx\exp(V)^{-1}\exp(U)$ (1)
 Section 3.1 will show that this is the key to being able to encode relative positions.
@@
 These form a special Lie group that is commutative, allowing us to strengthen the statement in equation (1) to
-$exp(x-y)=exp(x)exp(y)^{-1}=exp(y)^{-1}exp(x).$ (2)
+$\exp(U-V)=\exp(U)\exp(V)^{-1}\boldsymbol{=}\exp(U)^{-1}\exp(V).$ (2)
 Our work examines the tradeoff between using the stronger property in equation (2) or increased capacity and the weaker property (1).
-In short, we find that benefit of representational flexibility exceeds the benefit of exactly honoring equation (2).
 3 Methods
@@
 Before introducing the pseudocode we explain how LieRE modifies the attention mechanism and how this relates to Lie groups.
-We start by using a linear map to embed the positions as skew-symmetric matrices $P_{i}$ and $P_{j}$ and compute high dimensional rotations as $R_{i}=\exp(P_{i})$ and $R_{j}=\exp(P_{j})$.
-Recall that, for every pair of tokens, the attention mechanism computes the inner products between their key and query vectors, $k_{i}^{T}q_{j}$.
+We start by using a learnable linear map to embed the positions as skew-symmetric matrices $P_{i}$ and $P_{j}$ and compute high dimensional rotations as $R_{i}=\exp(P_{i})$ and $R_{j}=\exp(P_{j})$.
+Learning in the space of skew symmetric matrices allows us to sidestep some of the difficulty that would come from learning on the manifold of rotation matrices.
+Recall that, for every pair of tokens, the attention mechanism computes the inner products between their key and query vectors, $k_{i}^{T}q_{j}$, as shown in figure 1.
 We encode the positions by multiplying by the rotation matrices in the prior part.
@@
 In other words, the inner product automatically computes the relative position encoding.
-Note that this is exactly the mechanism behind the original RoPE paper.
+Note that this is the mechanism behind the original RoPE paper.
 The difference is that LieRE uses the full set of high dimensional rotations and relaxes the constraint that equation 1 holds with equality.
@@
 LieRE modifies the attention mechanism by applying a rotation to the keys and queries before computing their inner products.
-However, where RoPE applies a block-diagonal $n\times n$ with 2D rotation matrices as blocks, LieRE applies a $n\times n$ rotation matrix.
+However, where RoPE applies a block-diagonal rotation matrix with 2D rotation matrices as blocks, LieRE applies a $n\times n$ rotation matrix (Figure 1).
 Using the fact that every rotation matrix can be represented as the matrix exponential of a skew-symmetric matrix, we parametrize the rotations with generators before the matrix exponential.
@@
 $R_{LieRE}:=\exp(Ax).$
-Then, for each attention mechanism, the keys and queries for a particular token are updated as $\mathrm{key^{\prime}}:=R_{LieRE}\cdot\mathrm{key},\mathrm{query^{\prime}}:=R_{LieRE}\cdot\mathrm{query}$ We present attention with RoPE and LieRE side by side in algorithms 1 and 2.
-Algorithm 1 RoPE Attention (prior work)
+Then, for each attention mechanism, the keys and queries for a particular token are updated as $\mathrm{key^{\prime}}:=R_{LieRE}\cdot\mathrm{key},\mathrm{query^{\prime}}:=R_{LieRE}\cdot\mathrm{query}$ We present attention with LieRE and LieRE-Commute side by side in algorithms 1 and 2.
+Algorithm 1 RoPE Attention (prior work for text), rot=rotated
 1:procedure RoPE($X,p,d$)
@@
 3:     for $i\leftarrow 0$ to $d$ step $2$ do
+4:         $X_{\text{rot}}^{i}\leftarrow X^{i}\cos p\theta^{i}-X^{i+1}\sin p\theta^{i}$
+5:         $X_{\text{rot}}^{i+1}\leftarrow X^{i}\sin p\theta^{i}+X^{i+1}\cos p\theta^{i}$
+6:     end for
+7:     return $X_{\text{rotated}}$
+8:end procedure
+9:procedure RoPEAttention($Q,K,V$)
+10:     $p\leftarrow\text{tokenPositions}$
 […]
-5:         $X_{\text{rotated}}[i]\leftarrow X[i]\cdot\text{cos\_pos}_{i}-X[i+1]\cdot\text{sin\_pos}_{i}$
-6:         $X_{\text{rotated}}[i+1]\leftarrow X[i]\cdot\text{sin\_pos}_{i}+X[i+1]\cdot\text{cos\_pos}_{i}$
-7:     end for
-8:     return $X_{\text{rotated}}$
-9:end procedure
-10:procedure RoPEAttention($Q,K,V$)
-11:     $p\leftarrow\text{tokenPositions}$
-[…]
-13:     $K_{\text{rotated}}\leftarrow\textsc{RoPE}(K,p,d)$
-14:     $Q_{\text{rotated}}\leftarrow\textsc{RoPE}(Q,p,d)$
-15:     $\text{Attention}\leftarrow\text{softmax}(\frac{Q_{\text{rotated}}K_{\text{rotated}}^{T}}{\sqrt{d}})V$
-16:     return Attention
-17:end procedure
-Algorithm 2 LieRE Attention (our work)
+12:     $K_{\text{rot}}\leftarrow\textsc{RoPE}(K,p,d)$
+13:     $Q_{\text{rot}}\leftarrow\textsc{RoPE}(Q,p,d)$
+14:     $\text{Attention}\leftarrow\text{softmax}\left(\frac{Q_{\text{rot}}K_{\text{rot}}^{T}}{\sqrt{d}}\right)V$
+15:     return Attention
+16:end procedure
+Algorithm 2 LieRE Attention (our work), rot=rotated
 1:procedure LieRE($X,p,A$)
@@
 […]
-4:     return $X_{\text{rotated}}$
+4:     return $X_{\text{rot}}$
 5:end procedure
@@
 7:     $p\leftarrow\text{tokenPositions}$
-8:     $K_{\text{rotated}}\leftarrow\textsc{LieRE}(K,p,A)$
-9:     $Q_{\text{rotated}}\leftarrow\textsc{LieRE}(Q,p,A)$
-10:     $\text{Attention}\leftarrow\text{softmax}\left(\frac{Q_{\text{rotated}}K_{\text{rotated}}^{T}}{\sqrt{\text{dim}(K)}}\right)V$
+8:     $K_{\text{rot}}\leftarrow\textsc{LieRE}(K,p,A)$
+9:     $Q_{\text{rot}}\leftarrow\textsc{LieRE}(Q,p,A)$
+10:     $\text{Attention}\leftarrow\text{softmax}\left(\frac{Q_{\text{rot}}K_{\text{rot}}^{T}}{\sqrt{\text{dim}(K)}}\right)V$
 11:     return Attention
@@
 Outside of the generator scaling experiments, we set the rotation dimension equal to the head dimension and apply rotation on a per-attention-head basis.
-Note that if we are applying our method to images it is possible to impose a sparsity structure on the generator that recovers RoPE-mixed without the absolute position embeddings [heo2024rotary].
+Note that if we apply sparsity structure that enforces 2x2 blocks we recover commutativity.
+We refer to this configuration and LieRE-Commute.
+If we then have separate learnable parameters for every layer and head, we recover RoPE-mixed [heo2024rotary].
 4 Experiments
 4.1 Model and Training Parameters
-In order to isolate the benefit of LieRE, we use the standard transformer backbone modified to be able to switch between relative position encoding types.
-The backbone is configured as ViT-B, with 12 layers, a hidden dimension of 768, and an intermediate dimension of 3096.
-All experiments use dropout of 0.1.
+In order to isolate the effect of changing the position encoding, we use a standard [^1] https://github.com/kentaroy47/vision-transformers-cifar10 transformer backbone modified to be able to switch between relative position encoding types.
+We use the standard backbone sizes of ViT-Tiny, ViT-B and ViT-L [dosovitskiy2020image].
 All experiments use RandAugment [cubuk2020randaugment].
-We avoid using pre-trained weights in order to maximize the comparability of results between methods. All trainers use ADAM with the default parameters for pytorch.
-In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters with LieRE and use the same default hyperparameters for all experiments (Appendix A).
+We avoid using pre-trained weights in order to maximize the comparability of results between methods.
+In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters with LieRE and use the same default hyperparameters for all experiments (Appendix B).
 4.2 Datasets and Tasks
-Our experiments are designed to evaluate the efficacy of LieRE as a position encoding across a diverse set of modalities.
-We evaluate LieRE on the classification of 2D (images) and 3D (videos and medical images) data.
-For 3D data and ImageNet (2D), we focus on accuracy.
-For CIFAR100 (2D), where training is less resource intensive, we also evaluate LieRE’s data and training compute efficiency.
+Our experiments are designed to evaluate the efficacy of LieRE as a position encoding across both 2D and 3D data.
+We evaluate LieRE on the classification of 2D (images) and 3D (videos) data.
+For 3D data and ImageNet-1k (2D), we focus on accuracy.
+For CIFAR-100 (2D), where training is less resource intensive, we also evaluate LieRE’s data and training compute efficiency.
 4.2.1 2D Classification
-For 2D data we evaluate performance on the CIFAR100 and ImageNet image classification task [krizhevsky2009learning, deng2009imagenet].
-All hyperparameters and other details necessary for the reproducibility of experiments can be found in appendix A.
+For 2D data we evaluate performance on the CIFAR-100 and ImageNet-1k image classification task [krizhevsky2009learning, deng2009imagenet].
 We partition our evaluation of performance into four parts.
-In the first part, we examine accuracy across a range of model architectures on both CIFAR100 and ImageNet.
-To ensure a fair comparison, we implement all positional encodings on the same [standard] transformer backbone.
-These experiments are meant to identify the best-performing method of encoding positions.
-In the second part, we take advantage of the relatively modest amount of compute resources necessary to train a model for CIFAR 100 to examine LieREs impact on data efficiency.
-Here we limit our focus to LieRE, RoPE-Mixed, VisionLlama embeddings and absolute position embeddings.
-We train every model on 30%, 40%,…,90% subsets of CIFAR100 in order to evaluate the sensitivity of each model to the amount of training data.
-Since training starts to overfit quite early for the smaller datasets, we use the maximum validation accuracy for each run when comparing models.
+In the first part, we examine accuracy across a range of model architectures on both CIFAR-100 and ImageNet-1k and compare to absolute [dosovitskiy2020image] position encoding and recent related work, RoPE-Mixed [heo2024rotary], VisionLlama embeddings [chu2024visionllama].
+In the second part, we take advantage of the relatively modest amount of compute resources necessary to train a model for CIFAR-100 to examine LieREs impact on data efficiency.
 We also measure training compute efficiency by comparing the number of training steps necessary to achieve a fixed level of validation accuracy.
-In the third part, we evaluate the scaling laws of the LieRE generator capacity by imposing a block diagonal structure on the generator that allows us to vary the added capacity.
-With a head dimension of $64$ LieRE adds $4032$ parameters to a transformer designed for $2D$ data.
-While this is a small number of parameters, they lie in a critical point in the architecture, and understanding the marginal impact of this capacity can help inform computational tradeoffs.
-The block diagonal generator structure results in a block-diagonal rotation matrix after the matrix exponential.
-The parameters for each block are learned separately.
-This setup allows us to vary the capacity of LieRE without modifying the rest of the model.
-In the last part, we measure how much different models depend on the positional information in the image by shuffling the patches.
-A higher accuracy drop means the model relies more on the positions of the patches during inference.
+In the third part, we evaluate the impact of LieRE with various scales.
+We vary the capacity of the transformer backbone with corresponding to ViT-T, ViT-B and ViT-B, as proposed in [dosovitskiy2020image], the number of learned LieRE generators (one, per-attention-head and per-layer) and the LieRE generator capacity by imposing a block diagonal structure on the generator that allows us to vary the added capacity.
+In the fourth part, we measure how much different models depend on the positional information in the image/video by shuffling the patches.
+A higher accuracy drop with randomly shuffled patches means the model relies more on the positions of the patches during inference.
 4.2.2 3D Classification
 In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-mixed and absolute encoding similar to the previous section [arnab2021vivit, heo2024rotary].
-For the 3D experiments we examine video classification performance in the UCF101 dataset [soomro2012ucf101] and hemorrhage classification on the RSNA dataset [rsna2019data].
-We treat the RSNA dataset as a binary classification problem, asking the model to identify whether a 3D brain scan has a hemorrhage or not.
-Again, we did not optimize any hyperparameters for the LieRE model.
-We additionally use Mixup described standard used augmentation method for videos [zhang2017mixup].
-The full set of hyperparameters may be found in appendix A.
+For the 3D experiments, we examine video classification performance in the UCF101 dataset [soomro2012ucf101].
+Again, we did not optimize any hyperparameters for the LieRE model and used the dataloader from [tong2022videomae].
+The full set of hyperparameters may be found in appendix B.
 5 Results
 5.1 Accuracy
-For 2D images, we show that the LieRE-based transformer outperforms the DeiT by 5.5% [touvron2022deit], RoPE adaptation in VisionLlama [chu2024visionllama] by 3.6% and RoPE-Mixed by 2.7% on Cifar100 [heo2024rotary].
-For 3D inputs (video and 3D medical images), we observe an accuracy improvement of LieRE-based transformer over the Absolute Position Embedding of up to 6.7 % and RoPE-based transformers by up to 2.5% in accuracy.
-Table 1: 2D image classification on Cifar100 and ImageNet [krizhevsky2009learning, deng2009imagenet], equivalent to DeiT
-Method # params Cifar100 Accuracy (%) ImageNet Accuracy (%)
-Absolute Position E. 85.1M 63.9 66.1
-VisionLlama RoPE 85.1M 65.8 -
-RoPE-Mixed 85.1M 66.7 68.5
-LieRE (ours) 85.1M 69.4 68.8
-Table 2: 3D video classification on RSNA hemorrhage prediction and UCF101 dataset [soomro2012ucf101, rsna2019data], equivalent to Vivit (spatio-temporal)
-Method # params UCF101 Accuracy (%) RSNA Accuracy (%)
-Absolute Position E. 88.7M 44.4 80.7
-RoPE-Mixed 88.7M 48.6 81.9
-LieRE (ours) 88.7M 51.1 82.7
+For 2D images, we show that the LieRE-based transformer outperforms the DeiT by a relative performance improvement of 9.4% [touvron2022deit], RoPE adaptation in VisionLlama [chu2024visionllama] by 6.7% and RoPE-Mixed by 9.7% on CIFAR-100 [heo2024rotary].
+For 3D inputs (video), we observe an relative accuracy improvement of LieRE-based transformer over the absolute position embedding of up to 25.5 % and RoPE-inspired position encodings at least to 6.1% in accuracy.
+Table 1: 2D image and 3D video classification Top-1 Accuracy (95% confidence intervals) results.
+All models use 85.1M parameters for 2D tasks and 88.7M parameters for 3D task [krizhevsky2009learning, deng2009imagenet, soomro2012ucf101, rsna2019data] equivalent to DeiT, ∗∗ equivalent to Vivit (spatio-temporal).
+Method CIFAR-100 ImageNet-1k UCF101
+Abs. Pos. E.∗,∗∗ 63.9 (62.9-65.8) 66.1 (65.7-66.5) 42.7 (42.3-43.1)
+VisionLlama RoPE 65.5 (64.6-66.5) 65.4 (65.0-65.8) 50.5 (50.1-50.9)
+RoPE-Mixed 63.7 (61.8-65.6) 65.9 (65.5-66.3) 45.7 (45.2-46.0)
+LieRE-Commute 66.9 (66.0 - 67.9) 68.4 (67.9-68.9) 52.1 (51.7-52.5)
+LieRE 69.9 (68.9-70.8) 69.5 (69.1-69.9) 53.6 (53.2-54.0)
 5.2 Data efficiency
 We observe that transformers based on LieRE exhibit greater data efficiency compared to leading transformer architectures for 2D images on the CIFAR-100 dataset.
-For instance, using only 40% of the dataset, the LieRE-based transformer outperforms all comparable methods by 4.47%.
+For instance, when using 40% of the dataset, the LieRE-based transformer outperforms all comparable methods by 4.5 percentage points.
 Additionally, there is a mild trend indicating that the performance gap widens as the amount of data decreases (Figure 3(b)).
-(a) Performance scales with generator dimension (learnable parameters)
-(b) Data ablation on Cifar100
-Figure 2:
-5.3 Generator Dimension Scaling Laws
-LieRE adds a small amount of capacity to the model ($4032$ parameters in our case) but still represents the largest amount of learnable capacity around relative position encodings to date.
-This leads to the natural question of how helpful marginal capacity is.
-We vary capacity by enforcing a block diagonal structure on the generator. varying the block size allows us to approach "full" LieRE. Note that $2\times 2$ blocks correspond exactly to RoPE-Mixed.
+(a) Performance scales with generator dimension the number of LieRE parameters, compared to LieRE with one skew-symmetric matrix learned for the entire model (stem).
+(b) Data ablation on for different position embeddings on CIFAR-100.
+5.3 Model Scaling
+We examine three axes of capacity scaling: transformer backbone parameters, LieRE generator parameters and learning separate LieRE generators across heads and layers.
+Here we find a surprising result that RoPE-mixed can be substantially improved by sharing parameters across attention heads and layers and not learning them separately.
+5.3.1 Transformer backbone capacity
+We also evaluate how the inclusion of LieRE affects performance across model sizes on the CIFAR-100 in Figure 4(a).
+We observe that LieRE retains a statistically significant lead in performance across all three model sizes.
+Refer to caption (a) Performance behavior on CIFAR-100 (2D Image Classification) over ViT-Tiny (22M), ViT-Base (85M), ViT-Large (302M) for LieRE Rope-Mixed and Absolute Encoding (Appendix, table 5).
+Refer to caption (b) The LieRE spatial encoding allows the model to match the performance of absolute position encodings with substantially less training time.
+5.3.2 Generator parameters
+LieRE adds a small amount of capacity to the model ($4032$ parameters for the ViT-B backbone we use for most experiments), leading to the natural question of how helpful the marginal capacity is.
+We vary capacity by enforcing a block diagonal structure on the generator.
+varying the block size allows us to approach "full" LieRE.
+Recall that using $2\times 2$ blocks recovers exact commutativity and is referred to as LieRE-commute.
+This configuration substantially outperforms RoPE-Mixed, despite differing only in that it uses the same parameters across heads and layers and RoPE-Mixed learns separate parameters.
 In figure 3(a) we evaluate accuracy versus block dimension. We observe the steepest increase at the very start, though accuracy continues to grow as the generator dimension increases.
+5.3.3 Sharing LieRE Parameters Across Heads and Layers
+We also examine whether learning separate positional encoding per attention head and per layer improves performance on CIFAR-100 in table 2.
+We conduct this experiment for both RoPE-Mixed and LieRE and observe that in both cases it is best to learn one set of parameters across heads and layers.
+This is the mechanism by which we are able to show larger improvements from adopting RoPE-Mixed than the original paper [heo2024rotary].
+We name this transformed form of RoPE-Mixed LieRE-Commute, since it is equivalent to LieRE with generator dimension 2 (see Figure LABEL:fig:generator_dataset).
+Table 2: Layer, Head and Stem Learnable Rotary Embeddings on CIFAR-100, ViT-B, equivalent to LieRE-Commute
+FLOP Stem Heads Layers Rope-Mixed LieRE Reference
+5.684G 63.0 64.5 [heo2024rotary]
+5.684G 63.6 65.1
+5.613G 68.3 69.9
+5.613G 66.9 69.9 (ours)
 5.4 Compute efficiency
-Figure 3: The LieRE spatial encoding allows the model to match the performance of absolute position encodings with substantially less training time.
 Training transformers can necessitate substantial computational resources, which can hinder equitable access to research and development of machine learning methods.
-We demonstrate that the LieRE-based transformer requires 3.5 times less training time to achieve comparable performance to the Absolute Position Embedding baseline (as used in DeiT III [touvron2022deit]).
+We demonstrate that the LieRE-based transformer requires 3.5 times less training epochs on Cifar100 to achieve comparable performance to the Absolute Position Embedding baseline (as used in DeiT III [touvron2022deit]).
 This represents the largest reduction in training time compared to recent works such as VisionLlama and RoPE-Mixed [chu2024visionllama, heo2024rotary].
 We note that this is the first training efficiency comparison of these recent methods.
-Figure 3 shows the amount of allowable compute reduction to achieve the same accuracy achieved by absolute position encodings (DEIT baseline) after 200 epochs.
+Figure 4(b) shows the amount of allowable compute reduction to achieve the same accuracy achieved by absolute position encodings (DeiT baseline) after 200 epochs.
 LieRE demonstrates the largest win, allowing a 3.5X reduction in training compute to achieve the same accuracy.
-While all methods display better performance with more training compute, a $3.5x$ reduction allows for significantly faster experimentation and makes a fixed level of performance more accessible to those with limited access to compute resources.
+5.5 FLOP analysis
+We find that since all methods we examine introduce a computational cost that is at most linear in the number of tokens, and runtime is dominated by the quadratic attention component, there is no substantial difference in computational efficiency between the methods.
+We list inference FLOP of the various methods in table 3.
+Table 3: FLOP analysis with percentage increase compared to absolute position encodings
+Position Enc. ViT-Tiny (22M) ViT-Base (85M) ViT-Large (302M)
+Abs. Pos. E. 0.963G 5.607G 19.856G
+VisionLlaMA RoPE 0.963G (+0.001%) 5.607G (+0.002%) 19.856G (+0.000%)
+RoPE-Mixed 0.963G (+0.001%) 5.609G (+0.036%) 19.863G (+0.035%)
+LieRE-Commute (ours) 0.963G (+0.000%) 5.608G (+0.018%) 19.858G (+0.005%)
+LieRE (ours) 0.970G (+0.727%) 5.613G (+0.107%) 19.869G (+0.065%)
 5.6 Patch shuffling
@@
 We evaluate models using the decline in accuracy when evaluating on shuffled patches.
-We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE encodes meaningful relative positions between tokens.
-The complete results are displayed in table LABEL:tab:shuffling_results for cifar100 and table LABEL:tab:shuffling_results_3D for UCF101.
-Table 3: Relative drop in accuracy for 2D image classification after shuffling patch order, Cifar100
-Method before shuffling $\uparrow$ after shuffling $\downarrow$ Accuracy Drop (%) $\uparrow$
-Absolute Position E. 63.9 19.6 69.3
-VisionLlama RoPE 65.8 29.7 54.8
-RoPE-Mixed 66.7 16.5 75.4
-LieRE (ours) 69.4 15.9 77.1
-Table 4: Accuracy for Video recognition after patch shuffling, UCF101 dataset
-Method before shuffling $\uparrow$ after shuffling $\downarrow$ Accuracy Drop (%) $\uparrow$
-Absolute Position E. 44.4 44.0 0.1
-RoPE-Mixed 48.6 36.6 24.7
-LieRE (ours) 51.1 32.2 36.9
+We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE models are most capable at using positional information.
+The complete results are displayed for CIFAR-100 and table for UCF101 (Table 4) .
+Table 4: Relative accuracy drop for 2D image classification (CIFAR-100) and Video recognition (UCF101) after patch shuffling
+CIFAR-100 (2D) UCF101 (3D)
+Method Before After Drop(%) Before After Drop(%)
+[…]
+Abs. Pos. E. 63.9 19.6 69.3 42.7 42.5 0.0
+VisionLlama RoPE 65.8 29.7 54.8 50.5 36.2 28.5
+RoPE-Mixed 63.6 38.7 39.0 45.7 37.9 17.1
+LieRE-Commute 66.9 16.5 75.4 52.1 37.2 28.6
+LieRE 69.9 15.9 77.1 53.6 24.7 53.9
 6 Limitations
-While LieRE shows promising results across multiple modalities and input dimensionalities, there are a few limitations worth noting:
-1.
-Compatibility with architectures that do not use attention: Our method is specifically designed to modify the inner product, making it compatible with most attention schemes, including original attention and linear attention.
+While LieRE shows promising results across multiple modalities and input dimensionalities, there are a few limitations worth noting.
+Our method is specifically designed to modify the inner product, making it compatible with most attention schemes, including original attention and linear attention.
 However, this specificity may limit its applicability to other architectures, like convolutional neural networks, that do not rely on inner product-based attention mechanisms.
-Future work could explore adaptations of our method to a wider range of architectures.
-2.
-Limited to positions in $\mathbb{R}^{n}$: The current formulation of our method is designed to encode positions in $\mathbb{R}^{n}$.
-While this is sufficient for many applications, it may not be directly applicable to tasks that require encoding poses in $SE(3)$, like robotics.
+Future work could explore adaptations of our method to a wider range of architectures.Furthermore, the current formulation of our method is designed to encode vector positions in $\mathbb{R}^{d}$.
+While this is sufficient for many applications, it may not be directly applicable to tasks that require encoding poses in $SE(3)$ like robotics.
 Further research may be necessary to adapt our method to effectively handle such representational requirements.
@@
 7 Broader Impacts
-LieRE, our proposed method for encoding positional information in attention mechanisms, has demonstrated substantial improvements in 2D and 3D image classification tasks.
-We are particularly excited about how LieRE can expand the applicability of RoPE-based transformers to $n$ dimensions.
-This lays down the tracks to apply that same relative position within and across modalities and settings.
-1.
-One Backbone for All Modalities: LieRE enhances the generalizability of attention mechanisms, making them more versatile and adaptable to a wider range of modalities.
-Once the input is encoded as tokens and positions, no additional changes are necessary in order to support a new modality.
-By providing a flexible and efficient way to encode positional information, LieRE can facilitate the application of attention mechanisms to diverse data types, such as audio, video, multi-dimensional sensor data or temporal data with diverse time intervals.
-This increased generalizability can lead to more effective transfer learning and cross-pollination of ideas between different research areas, accelerating progress and innovation in the field.
-2.
-Improving accessibility of attention-based architectures for low-data and low-compute regimes: This is particularly valuable for applications where labeled spatial data is scarce or expensive to acquire, such as in medical imaging and robotics.
-By lowering the barriers to entry, LieRE can enable the application of attention mechanisms to a broader set of problems, including those in resource-constrained settings or niche domains.
-3.
-Enabling more efficient and sustainable AI: The reduced training costs and improved data efficiency associated with LieRE can contribute to the development of more efficient and sustainable AI systems.
-By requiring fewer computational resources and less training data, LieRE can help mitigate the environmental impact of large-scale model training.
-This is particularly important as the field of AI continues to grow and the demand for computational resources increases.
-In conclusion, LieRE has the potential to expand the generalizability of attention mechanisms, improve their accessibility for spatial data, and contribute to the development of more efficient and sustainable AI systems.
+LieRE, our proposed method for encoding positional information in attention mechanisms, has demonstrated improvements in 2D image and video classification tasks.
+We are particularly excited about how LieRE can expand the applicability transformers to $n$ dimensional inputs.
+This may lay down the tracks to apply that same relative position encoding within and across modalities and settings.
 8 Conclusion
 In this paper, we introduced Lie group Relative position Encodings (LieRE), a position encoding that can effectively encode relative position information for attention mechanisms across modalities and input dimensionalities.
-Through extensive experiments on 2D image classification (CIFAR100, ImageNet) and 3D medical image classification (RSNA brain CT hemorrhage detection), we demonstrated that LieRE achieves superior performance compared to existing positional encoding methods.
+Through experiments on 2D image classification (CIFAR-100, ImageNet-1k) and 3D video classification (UCF101), we demonstrated that LieRE achieves better performance compared to existing positional encoding methods.
 Beyond improving accuracy, LieRE also exhibits data and compute efficiency.
-On CIFAR100, LieRE requires 3.5 times less training computed to match the performance of the baseline model with absolute position encodings.
+On CIFAR-100, LieRE requires 3.5 times less training computed to match the performance of the baseline model with absolute position encodings.
 Furthermore, LieRE can outperform the baseline trained on the full dataset while using only 70% of the training data, highlighting its data efficiency.
 The key advantages of LieRE include its simplicity, flexibility, and ease of adaptation to new modalities.
-By requiring no changes to the tokenizer other than outputting positions in $\mathbb{R}^{n}$ and no other code changes, LieRE provides a unified and efficient approach for transformers to process and learn from various data modalities within a single architecture.
-We hope that this empowers the field in building models that can understand concepts.
-Funding and Acknowledgments We would like to thank Maksim Maydanskiy for helping us understand Lie groups and Lie group representations.
-We would like to thank Aradhana Sinha for suggesting the shuffling experiment and providing early feedback on the paper.
-Sophie Ostmeier was supported by the German Research Foundation (DFG), Walter-Benjamin fellowship (ID: 517316550).
-This project was supported by Google Cloud and Azure credits.
+By requiring no changes to the tokenizer other than outputting positions and no other code changes, LieRE may provide a unified and efficient approach for transformers to process and learn from various data modalities within a single architecture.
 […]
-Appendix A Experiment Hyperparameters
-The backbone for all experiments is configured as ViT-B, with 12 layers, a hidden dimension of 768, and an intermediate dimension of 3096. We use dropout of 0.1.
-A.1 2D Image Classification
+Appendix A Appendix
+Appendix B Experiment Hyperparameters
+The backbone for all experiments is configured as ViT-B, with 12 layers, a hidden dimension of 768, and an intermediate dimension of 3096.
+We use a dropout of 0.1.
+We used CLS pooling in our implementation to facilitate comparability with existing literature in the field.
+Further experiments revealed substantial performance improvement with mean pooling and LieRE.
+We use the pytorch lightning framework for all experiments [falcon2019pytorch].
+B.1 2D Image Classification
 The CIFAR experiments where trained on 8xL4 GPUs with 24GB of VRAM each and all took under 30 minutes to complete.
@@
 We use an effective batch size of $512$.
-We use a patch size of $4\times 4$ on the original $32\times 32$ image for Cifar100 and a patch size of $16\times 16$ on the randomly cropped and resized $224\times 224$ image.
+We use a patch size of $4\times 4$ on the original $32\times 32$ image for CIFAR-100 and a patch size of $16\times 16$ on the randomly cropped and resized $224\times 224$ image.
 All vision experiments used RandAugment [cubuk2020randaugment].
@@
 The hyperparameters were tuned with RoPE-mixed and selected before conducting the LieRE trainers as to ensure a fair comparison.
-A.2 3D Medical Image and Video Classifications
+B.2 3D Video Classifications
 The 3D classification experiments were conducted on either $1008\times A$ 100 40GB GPUs or $1004\times A$ 100 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size.
-Similar to 2D classification, we use an initial learning rate of $41E-$ 4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $4\times 16\times 16$ on the randomly cropped and resized $32\times 224\times 224$ video/image.
+Similar to 2D classification, we use an initial learning rate of $41E-$ 4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $2\times 16\times 16$ on the randomly cropped and resized $32\times 224\times 224$ video/image.
 We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$.
-We use Mixup for all 3D experiments [tong2022videomae, zhang2017mixup].
+B.3 Scaling Model Size
+Table 5: Comparison of Position Encoding Methods for Different ViT Models, Accuracy (bootstrapped 95%CI)
+Position Encoding ViT-Tiny ViT-Base ViT-Large
+Abs. Pos. E. 57.2 (56.2 - 58.1) 63.9 (62.9 - 65.8) 60.5 (59.5 - 61.4)
+VisionLlaMA RoPE 58.2 (57.2 - 59.2) 65.5 (64.6 - 66.5) 62.3 (60.4 - 64.2)
+RoPE-Mixed 59.6 (58.7 - 60.6) 63.7 (61.8 - 65.6) 65.0 (64.0 - 65.9)
+LieRE-Commute 63.2 (62.2 - 64.1) 66.9 (66.0 - 67.9) 65.1 (64.1 - 66.0)
+LieRE 64.7 (63.8 - 65.7) 69.9 (68.9 - 70.8) 69.2 (68.3 - 70.1)
+B.4 Python implementation of skew-symmetric matrix
+1generator_raw_params = nn.Parameter(
+2 torch.rand(
+3 input_dimensionality,
+4 head_dim,
+5 head_dim,
+6 ) * 2 * math.pi # optional, inspired from RoPE-Mixed paper
+7)
+8
+9upper_triangle = (
+10 torch.triu(generator_raw_params, diagonal=1)
+11)
+12skew_bases = upper_triangle - torch.transpose(upper_triangle, -1, -2)
+13in_basis_positions = (
+14 positions.reshape(list(positions.shape) + [1] * 2) * skew_bases
+15)
+16generator_pos = torch.sum(in_basis_positions, dim=-3)
+17rotation = torch.matrix_exp(generator_pos.to(dtype=torch.float32)).to(dtype=positions.dtype)

### v3

@@
-custom.bib
+LieRE: Lie Rotational Positional Encodings
 Abstract
-While Rotary Position Embeddings (RoPE) for large language models have become widely adopted, their application for other modalities has been slower.
-Here, we introduce Lie group Relative position Encodings (LieRE) that goes beyond RoPE in supporting n-dimensional inputs.
-We evaluate the performance of LieRE on 2D and 3D image classification tasks and observe that LieRE leads to marked relative improvements in performance (up to 9.7% for 2D and up to 25.5% for 3D), training efficiency (3.5x reduction), data efficiency (30%) compared to the baselines of DeiT III, RoPE-Mixed and Vision-Llama.
-https://github.com/Stanford-AIMI/LieRE
+Transformer architectures rely on position encodings to capture token dependencies.
+Rotary Position Encoding (RoPE) has emerged as a popular choice in language models due to its efficient encoding of relative position information through key-query rotations.
+However, RoPE faces significant limitations beyond language processing: it is constrained to one-dimensional sequence data and, even with learnable phases, offers limited representational capacity.
+We address these challenges with Lie Relative Encodings (LieRE), which replaces RoPE’s block-2D rotation matrix with a learned, dense, high-dimensional rotation matrix of variable density.
+Through extensive evaluation on three image datasets across 2D and 3D classification tasks, LieRE achieves 2% relative improvement over state-of-the-art baselines on 2D tasks and 1.5% on 3D tasks, while demonstrating superior generalization to higher resolutions.
+Our implementation is computationally efficient, with results reproducible on 4 A100 GPUs in 30 minutes on CIFAR100, and we release our code to facilitate further research.
+Machine Learning, ICML
 1 Introduction
-While the attention mechanism has achieved widespread use, especially as part of the transformer architecture, attention is invariant to the order of its inputs and requires another mechanism to capture positional information of input tokens [vaswani2017attention].
-This has spurred a line of work in the subarea of positional encodings—methods of encoding positional information in attention mechanisms.
-In particular, Rotary Position Encoding (RoPE) has emerged as a technique for encoding relative positional information of text tokens in transformer-based models [RoFormer2024Su].
-RoPE’s ability to capture relative position information has made it a popular choice for open-source language foundation models such as LLaMA.
-In particular, RoPE implicitly captures relative positions.
-When the token in position $p_{1}$ attends to a token in position $p_{2}$, the effect of RoPE depends on $p_{1}-p_{2}$.
-Despite the success of RoPE in sequence tasks [touvron2023llama, chowdhery2023palm], it is designed for one-dimensional sequence data.
-This has resulted in slow adoption for modalities with higher dimensional data, such or data that includes a temporal dimension like videos.
-Our work aims to answer whether a single position encoding scheme can work well across both 2D and 3D modalities.
-If possible, this would enable the use of a simpler common model backbone across a variety of tasks.
-[…]
-Figure 1: LieRE sketch, where $A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x\;y\;z])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the first patch in the top left corner of the input with the position $P_{i}=(x,y,z)\in\mathbb{R}^{N}$.
+The attention mechanism, particularly within transformer architectures, has revolutionized machine learning across diverse domains.
+However, attention is inherently permutation—invariant—it cannot leverage the sequential order of its inputs as information.
+This fundamental limitation necessitates additional mechanisms to encode positional information, enabling models to capture sequential and spatial dependencies crucial for tasks ranging from natural language processing to image understanding (Vaswani et al., 2017).
+This challenge has sparked extensive research into positional encodings, which inject order information into the otherwise order-agnostic attention mechanism.
+The field has evolved from early approaches using fixed sinusoidal positional embeddings (Vaswani et al., 2017) to more sophisticated learned embeddings (Shaw et al., 2018a; Devlin et al., 2019; Dosovitskiy et al., 2020).
+Recent advances have introduced increasingly dynamic methods, including relative position representations (Shaw et al., 2018b; Dosovitskiy et al., 2020) and rotary positional embeddings (Su et al., 2024; Heo et al., 2024), demonstrating the critical role of position encoding in enabling transformers to effectively process ordered sequences.
+In particular, Rotary Position Encoding (RoPE) has emerged as an elegant solution for encoding relative positional information in transformers (Su et al., 2024).
+RoPE works by applying a rotation matrix to each token’s keys and queries, where the rotation angle depends on the token’s absolute position in the sequence.
+The key insight is that when two tokens interact through the attention inner product, their rotated representations naturally encode their relative distance.
+For example, when a token at position five attends to a token at position two, RoPE enables the model to understand that these tokens are three positions apart, regardless of their absolute positions in the sequence.
+This translation-invariant property makes RoPE particularly efficient at capturing position-dependent patterns in text, which has led to its adoption in popular open-source language models such as LLaMA and Mixtral.
+Figure 1: Overview of LieRE, where $A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x\;y\;z])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the first patch in the top left corner of the input with the position $P_{i}=(x,y,z)\in\mathbb{R}^{N}$.
 $d$ is the head dimension.
 The upper triangle contains learnable $\textrm{LieRE}_{\theta}$ parameter, while the lower triangle contains their negatives, -$\textrm{LieRE}_{\theta}$, reflecting the skew-symmetric nature of the matrix.
-[…]
-Figure 2: Comparisons of basis matrix $A$ of our work, Rope-Mixed [heo2024rotary], VisionLlama [chu2024visionllama].
+Despite RoPE’s success in sequential tasks (Touvron et al., 2023; Chowdhery et al., 2023), it faces several fundamental limitations.
+First, RoPE is designed for one-dimensional sequence data with an exponential decay in frequencies, making it suboptimal for spatial relationships in images or spatiotemporal patterns in videos where distant pixels may have strong correlations.
+Second, when handling higher dimensional data, the challenge compounds significantly–learned relative position encodings must capture exponentially many relative positions for $n$-dimensional data, as demonstrated in 2D image data where positions must be encoded both horizontally and vertically (Shaw et al., 2018a).
+Third, RoPE’s reliance on sparse, block-diagonal rotation matrices with handcrafted basis functions constrains its ability to learn complex spatial dependencies (Chu et al., 2024).
+Our key insight is that Lie group theory provides a natural framework for generalizing relative position encodings to higher dimensions.
+We introduce Lie Relative Encodings (LieRE), which replaces RoPE’s rotation matrices with learned, dense rotations derived from Lie group generators.
+LieRE learns a basis of skew-symmetric matrices ${A_{i}}$, computes rotation matrices $R(p)=\exp(\sum_{i=0}^{n}p_{i}A_{i})$ for $n$-dimensional positions, and applies these rotations to keys and queries in the attention mechanism.
+The relative positions are then naturally captured through the inner product of the rotated keys and queries.
+This approach addresses RoPE’s limitations in two key ways: (1) it handles higher-dimensional spatial relationships through Lie groups and (2) it increases representational capacity through variably dense, learned rotation matrices while requiring only 0.68% of parameters in a ViT-B model.
+Importantly, LieRE can be implemented with a single modification to existing architectures.
+To assess the impact of LieRE and other positional encodings on ViT performance, we evaluate several encoding schemes across diverse tasks, including 2D and 3D image classification.
+Additionally, we investigate a fundamental spatial reasoning task where models must identify arrow directions.
+Despite their sophistication, contemporary multimodal LLMs struggle with this seemingly simple task.
+Our experiments reveal that successful completion of this task specifically requires relative position encodings, highlighting their crucial role in spatial understanding.
+Figure 2: Comparisons of basis matrix $A$ of our work, Rope-Mixed (Heo et al., 2024), VisionLlama (Chu et al., 2024).
 First row: learnable parameter (yellow), not learnable parameter (blue) in basis matrix.
@@
 MHA := Multi-Head-Attention, FF := Feed Forward, L[index] := Layer index, RxH := Rotation Matrix per Head, $\textrm{LieRE}_{\theta}$ := the number of LieRE parameters.
-1.1 Contributions
-We introduce Lie Relative Encodings (LieRE), a mechanism that allows the attention mechanism to learn how to utilize relative spatial information of its inputs.
-We show that LieRE is effective on both 2D and 3D inputs of various modalities.
-Beyond improving classification accuracy, LieRE also reduces the amount of compute and data required during training to achieve a fixed accuracy.
-On the CIFAR100 task, this translates to 3.5x fewer training steps to achieve the same accuracy as the DeiT baseline and outperforming the DeiT baseline trained on the full set with only a 70% subset of the data.
-Furthermore, LieRE is simple to implement and adaptable to modalities, requiring only a tokenizer that also outputs a position in $\mathbb{R}^{d}$ in addition to a standard embedding.
-In order to aid the reproducibility of our results we will post our code on github.
-1.2 Related Work
-1.2.1 Position Encodings
-The fact that the original attention mechanism is invariant to the order of tokens has motivated the ongoing development of methods to incorporate positional information into the transformer architecture.
-We split our literature review into three broad classes of positional encodings: 1) absolute, 2) relative, and 3) contextual.
-Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the intput.
-Methods such as sinusoidal and learned absolute encodings add vectors to the input token embedding [vaswani2017attention, devlin2019bert, dosovitskiy2020image].
-Absolute position measures position with respect to an absolute reference, such as the start of the text or the top left corner of an image.
-Relative position encodings instead encode the relative positions of two tokens.
-One strategy is to learn an embedding for position deltas which can be incorporated into the attention mechanism [shaw2018self, liu2021swin, liu2022swinv2].
-However, this incurs quadratic computational cost in terms of the number of tokens.
+2 Related Work
+2.1 Position Encodings
+We organize our literature review into three broad classes of positional encodings: 1) absolute, 2) relative, and 3) contextual.
+Absolute position in this context refers to a position with respect to a consistent reference, usually the start of a tex or top left corner of an image.
+Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the input.
+Methods such as sinusoidal and learned absolute encodings directly add vectors to the input token embedding (Vaswani et al., 2017; Devlin et al., 2019; Dosovitskiy et al., 2020).
+Relative position encodings in contrast encode the relative positions of two tokens.
+One strategy is to learn an set of embeddings for position deltas which can be incorporated into the attention mechanism (Shaw et al., 2018b; Liu et al., 2021, 2022).
+However, this incurs quadratic computational cost in terms of the number of tokens as a separate embedding is required for every pair of tokens.
 Rotary Position Encodings (RoPE) avoid this cost by rotating the key and query vectors before the attention inner product.
-The algebraic properties of the block-diagonal rotation matrices used in RoPE ensures only relative positional information is captured in the attention mechanism [RoFormer2024Su].
-RoPE is quite widely used in open source LLMs including the PaLM, Llama and Mixtral models [touvron2023llama, chowdhery2023palm, jiang2024mixtral].
+The algebraic properties of the block-diagonal rotation matrices used in RoPE ensures only relative positional information is captured in the attention mechanism (Su et al., 2024).
+RoPE is quite widely used in open source LLMs including the PaLM, Llama and Mixtral models (Touvron et al., 2023; Chowdhery et al., 2023; Jiang et al., 2024).
 However, RoPE can perform poorly on inference for larger context sizes than the model was trained on.
 This has spurred an active line of work extending RoPE to longer contexts, work which we review later.
-We refer to the last category of positional as contextual position embeddings.
+We refer to the last category of positional encodings as contextual position embeddings.
 This category is defined by encodings that aim to capture semantic positional information lost in traditional absolute and relative position encodings, often motivated by reasoning or mathematical tasks.
-Contextual Position encodings achieve (CoPE) this by allowing the model to learn how the position is computed [golovneva2024contextual].
-Abacus embeddings enable transformers to learn how to handle arithmetic by better exposing the digit structure of numbers [mcleish2024transformers].
-1.2.2 Extensions of RoPE
-The efficiency and popularity of RoPE have led to several lines of work looking to extend it focused on specific domains.
+Contextual Position encodings achieve (CoPE) this by allowing the model to learn how the position is computed (Golovneva et al., 2024).
+Abacus embeddings enable transformers to learn how to handle arithmetic by better exposing the digit structure of numbers (McLeish et al., 2024).
+2.2 Extensions of RoPE
+The efficiency and popularity of RoPE have led to several lines of work building off of it.
 One notable one is context extension, which aims to address the fact that RoPE NLP models trained on short documents tend to perform poorly on long documents.
-Methods like NTK-aware context extension, YaRN and LongRoPE focus on enabling already trained models to handle long context, both with and without finetuning [ding2024longrope, peng2023yarn, tworkowski2024focused, chen2023extending].
-Another line of work has been specifically focused on adapting RoPE to image tasks.
-Both VisionLlama and RoPE-Mixed present relative position encodings inspired by RoPE that are able to encode 2D positional data [chu2024visionllama, heo2024rotary].
+Methods like NTK-aware context extension, YaRN and LongRoPE focus on enabling already trained models to handle long context, both with and without finetuning (Ding et al., 2024; Peng et al., 2023; Tworkowski et al., 2024; Chen et al., 2023).
+The final, and most relevant, line of work has been specifically focused on adapting RoPE to image tasks.
+Both VisionLlama and RoPE-Mixed present relative position encodings inspired by RoPE capable of encoding 2D positional data (Chu et al., 2024; Heo et al., 2024).
 The primary difference is that RoPE-Mixed has a learnable component, whereas VisionLlama does not.
@@
 In this section, we aim to provide a minimal introduction to Lie groups so that the reader is able to understand the mathematical motivations behind LieRE.
-Lie groups are well studied, especially in the context of representation theory, and standard texts including [fulton2013representation] are able to provide a more extensive introduction to the subject.
+Lie groups are well studied, especially in the context of representation theory, and standard texts including (Fulton & Harris, 2013) are able to provide a more extensive introduction to the subject.
 In this context, Lie groups are smooth sets of matrices that are closed under matrix multiplication and inversion.
-For every Lie group, the matrix exponential provides a smooth bijective map from a subset of $\mathbb{R}^{n\times n}$ (hereto referred to as the generator set) to the Lie group.
-The exponential map is a diffeomorphism and has the following key property for generators $U,V\in\mathbb{R}^{n\times n}$ close together:
+For every Lie group, the matrix exponential provides a smooth bijective map from a subset of $\mathbb{R}^{n\times n}$, also known as the Lie Algebra, to the Lie group.
+The exponential map is a diffeomorphism and has the following key property for $U,V\in\mathbb{R}^{n\times n}$ close together:
 $\exp(U-V)=\exp(-V+U)\approx\exp(V)^{-1}\exp(U)$ (1)
-Section 3.1 will show that this is the key to being able to encode relative positions.
-Both RoPE (in the context of text) and RoPE-Mixed use block-diagonal rotation matrices with 2D rotations as blocks.
-These form a special Lie group that is commutative, allowing us to strengthen the statement in equation (1) to
-$\exp(U-V)=\exp(U)\exp(V)^{-1}\boldsymbol{=}\exp(U)^{-1}\exp(V).$ (2)
-Our work examines the tradeoff between using the stronger property in equation (2) or increased capacity and the weaker property (1).
-3 Methods
-3.1 Rotations for Relative Positions in Attention
-Before introducing the pseudocode we explain how LieRE modifies the attention mechanism and how this relates to Lie groups.
-We start by using a learnable linear map to embed the positions as skew-symmetric matrices $P_{i}$ and $P_{j}$ and compute high dimensional rotations as $R_{i}=\exp(P_{i})$ and $R_{j}=\exp(P_{j})$.
-Learning in the space of skew symmetric matrices allows us to sidestep some of the difficulty that would come from learning on the manifold of rotation matrices.
-Recall that, for every pair of tokens, the attention mechanism computes the inner products between their key and query vectors, $k_{i}^{T}q_{j}$, as shown in figure 1.
-We encode the positions by multiplying by the rotation matrices in the prior part.
-In particular, we update the keys and queries as $k_{i}^{\prime}=k_{i}R_{i}$ and $q_{j}=q_{j}R_{j}$.
-This results in an updated inner product of $(R_{i}k_{i})^{T}R_{j}q_{j}=k_{i}R_{i}^{T}R_{j}q_{j}=k_{i}R_{i}^{-1}R_{j}q_{j}.$
-Recall that by equation 1, $R_{i}^{-1}R_{j}=\exp(P_{i})^{-1}\exp(P_{j})\approx\exp(P_{j}-P_{i})$.
-In other words, the inner product automatically computes the relative position encoding.
-Note that this is the mechanism behind the original RoPE paper.
-The difference is that LieRE uses the full set of high dimensional rotations and relaxes the constraint that equation 1 holds with equality.
-We make the description of the method more precise in the next section.
-3.2 LieRE
-LieRE modifies the attention mechanism by applying a rotation to the keys and queries before computing their inner products.
-However, where RoPE applies a block-diagonal rotation matrix with 2D rotation matrices as blocks, LieRE applies a $n\times n$ rotation matrix (Figure 1).
-Using the fact that every rotation matrix can be represented as the matrix exponential of a skew-symmetric matrix, we parametrize the rotations with generators before the matrix exponential.
-This skew-symmetric generator is computed via a learned linear map from the positions to the generator space.
-Let $x\in\mathbb{R}^{n}$ indicate the position of the token, and $A:\mathbb{R}^{d}\rightarrow\text{Skew}_{n}(\mathbb{R})$ be a learnable linear map to the set of skew-symmetric matrices.
-At the start of every forward pass, we compute the LieRE rotation matrix for each token as
-$R_{LieRE}:=\exp(Ax).$
-Then, for each attention mechanism, the keys and queries for a particular token are updated as $\mathrm{key^{\prime}}:=R_{LieRE}\cdot\mathrm{key},\mathrm{query^{\prime}}:=R_{LieRE}\cdot\mathrm{query}$ We present attention with LieRE and LieRE-Commute side by side in algorithms 1 and 2.
-Algorithm 1 RoPE Attention (prior work for text), rot=rotated
-1:procedure RoPE($X,p,d$)
-2:     $\theta\leftarrow\frac{1}{10000^{2i/d}}\quad\forall i\in[0,d)$
-3:     for $i\leftarrow 0$ to $d$ step $2$ do
-4:         $X_{\text{rot}}^{i}\leftarrow X^{i}\cos p\theta^{i}-X^{i+1}\sin p\theta^{i}$
-5:         $X_{\text{rot}}^{i+1}\leftarrow X^{i}\sin p\theta^{i}+X^{i+1}\cos p\theta^{i}$
-6:     end for
-7:     return $X_{\text{rotated}}$
-8:end procedure
-9:procedure RoPEAttention($Q,K,V$)
-10:     $p\leftarrow\text{tokenPositions}$
-[…]
-12:     $K_{\text{rot}}\leftarrow\textsc{RoPE}(K,p,d)$
-13:     $Q_{\text{rot}}\leftarrow\textsc{RoPE}(Q,p,d)$
-14:     $\text{Attention}\leftarrow\text{softmax}\left(\frac{Q_{\text{rot}}K_{\text{rot}}^{T}}{\sqrt{d}}\right)V$
-15:     return Attention
-16:end procedure
-Algorithm 2 LieRE Attention (our work), rot=rotated
-1:procedure LieRE($X,p,A$)
-[…]
-[…]
-4:     return $X_{\text{rot}}$
-5:end procedure
-6:procedure LieREAttention($Q,K,V,A$)
-7:     $p\leftarrow\text{tokenPositions}$
-8:     $K_{\text{rot}}\leftarrow\textsc{LieRE}(K,p,A)$
-9:     $Q_{\text{rot}}\leftarrow\textsc{LieRE}(Q,p,A)$
-10:     $\text{Attention}\leftarrow\text{softmax}\left(\frac{Q_{\text{rot}}K_{\text{rot}}^{T}}{\sqrt{\text{dim}(K)}}\right)V$
-11:     return Attention
-12:end procedure
-Outside of the generator scaling experiments, we set the rotation dimension equal to the head dimension and apply rotation on a per-attention-head basis.
-Note that if we apply sparsity structure that enforces 2x2 blocks we recover commutativity.
-We refer to this configuration and LieRE-Commute.
-If we then have separate learnable parameters for every layer and head, we recover RoPE-mixed [heo2024rotary].
+Both RoPE (in the context of text) and RoPE-Mixed use block-diagonal rotation matrices with 2D rotations as blocks 4.
+These form a special Lie group that is commutative, allowing us to strengthen the statement in (1) to
+$\exp(U-V)=\exp(U)\exp(V)^{-1}\boldsymbol{=}\exp(V)^{-1}\exp(U).$ (2)
+Our work examines the tradeoff between using the stronger property in (2) or increased capacity and the weaker property (1).
+Figure 3: Pseudocode for the Lie Rotary Embedding (LieRE) attention mechanism. Unlike RoPE, LieRE allows for dense rotation matrices.
+3.2 Attention Mechanism
+LieRE is a modification of the standard attention mechanism to introduce positional information, which we review below.
+The modification we propose is independent of whether we use multiple heads, so we focus on single-headed attention for simplicity.
+Let $X\in\mathbb{R}^{n\times d}$ be the set of input embeddings and $W_{Q},W_{K},W_{v}$ be learnable matrices.
+Let $Q=XW_{Q},K=XW_{K},V=XW_{V}$ be the keys, queries and values respectively.
+The outputs are computed as $\text{scores}=\frac{QK^{\top}}{\sqrt{d_{k}}}$, $\mathcal{W}=\text{softmax}(\text{scores})$ and final outputs $z=\mathcal{W}V$.
+We let $Q_{i}$ and $K_{i}$ denote the $i$th rows of $Q$ and $K$ respectively.
+4 Method
+LieRE is a simple modification to the attention mechanism that is presented in algorithm 3.
+Recall that we assume that positions are $n$-dimensional vectors, a matrix $A$ is skew-symmetric if $A^{T}=-A$, and that the matrix exponential of a skew-symmetric matrix, call it $A$, is always a high dimensional rotation matrix.
+When encoding positions $p\in\mathbb{R}^{n}$, LieRE learns a skew-symmetric basis of matrices $\{A_{i}\}$ for $i\in[n]$.
+It encodes a position by writing it in this basis, $\sum\limits_{i=0}^{n}p_{i}A_{i}$.
+We then map the resulting skew-symmetric matrix to a high-dimensional rotation via the matrix exponential.
+$R(p)=\exp\left(\sum\limits_{i=0}^{n}p_{i}A_{i}\right)$ (Figure 1).
+Learning in the space of skew-symmetric matrices allows us to sidestep some of the difficulty that would come from learning on the manifold of rotation matrices (Figure 2).
+Figure 4: Pseudocode for the Rotary Position Embedding (RoPE) attention mechanism.
+LieRE uses the rotation matrix computed above to modify the keys and queries of the standard attention mechanism.
+LieRE’s final step is to modify token $i$’s query and keys as $Q^{\prime}_{i}=R(p_{i})Q_{i}$ and $K^{\prime}_{i}=R(p_{i})K_{i}$.
+This modifies the score between tokens $i,j$ to be $X_{i}^{T}W_{Q}^{T}R(p_{i})^{T}R(p_{j})W_{K}X_{j}$.
+Recalling that $R^{T}=R^{-1}$ for any orthogonal matrix $R$ helps illustrate the encoding of relative positions in (1).
+Note that the only difference between LieRE and RoPE-Mixed is that the latter constrains the rotations to be block-diagonal with block size two.
+We include the pseudocode for the LieRE attention in algorithm 3 in addition to the standard RoPE attention (algorithm 4).
+In practice, we compute the rotation matrices at the start of the forward pass.
+By default, the skew symmetric bases are learned separately for every layer and attention head except in the experimental section focused on sharing parameters across heads and layers.
+Adjusting the skew-symmetric basis matrices’ block width allows us to incrementally adjust the capacity allocated towards position encoding.
+We specify the basis block width as a subscript, eg.
+LieRE8.
+When not specified, the block size is equal to the head dimension.
+If we set the block size to 2, we recover RoPE-Mixed (Heo et al., 2024).
+For a ViT-B model the head dimension is 64.
 5 Experiments
-4.1 Model and Training Parameters
-In order to isolate the effect of changing the position encoding, we use a standard [^1] https://github.com/kentaroy47/vision-transformers-cifar10 transformer backbone modified to be able to switch between relative position encoding types.
-We use the standard backbone sizes of ViT-Tiny, ViT-B and ViT-L [dosovitskiy2020image].
-All experiments use RandAugment [cubuk2020randaugment].
+In order to isolate the effect of changing the position encoding, we use a plain standard transformer backbone and training regime [^1] https://github.com/kentaroy47/vision-transformers-cifar10 with comparable baseline performance on CIFAR100 and ImageNet with (Lee et al., 2022).
+We implement relative position encoding types including LieRE, Rope-Mixed, Visionllama, and absolute position encodings.
+We use the standard backbone sizes of ViT-Tiny, ViT-B and ViT-L (Dosovitskiy et al., 2020).
+All experiments use RandAugment (Cubuk et al., 2020).
 We avoid using pre-trained weights in order to maximize the comparability of results between methods.
-In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters with LieRE and use the same default hyperparameters for all experiments (Appendix B).
-4.2 Datasets and Tasks
-Our experiments are designed to evaluate the efficacy of LieRE as a position encoding across both 2D and 3D data.
-We evaluate LieRE on the classification of 2D (images) and 3D (videos) data.
-For 3D data and ImageNet-1k (2D), we focus on accuracy.
-For CIFAR-100 (2D), where training is less resource intensive, we also evaluate LieRE’s data and training compute efficiency.
-4.2.1 2D Classification
-For 2D data we evaluate performance on the CIFAR-100 and ImageNet-1k image classification task [krizhevsky2009learning, deng2009imagenet].
-We partition our evaluation of performance into four parts.
-In the first part, we examine accuracy across a range of model architectures on both CIFAR-100 and ImageNet-1k and compare to absolute [dosovitskiy2020image] position encoding and recent related work, RoPE-Mixed [heo2024rotary], VisionLlama embeddings [chu2024visionllama].
-In the second part, we take advantage of the relatively modest amount of compute resources necessary to train a model for CIFAR-100 to examine LieREs impact on data efficiency.
-We also measure training compute efficiency by comparing the number of training steps necessary to achieve a fixed level of validation accuracy.
-In the third part, we evaluate the impact of LieRE with various scales.
-We vary the capacity of the transformer backbone with corresponding to ViT-T, ViT-B and ViT-B, as proposed in [dosovitskiy2020image], the number of learned LieRE generators (one, per-attention-head and per-layer) and the LieRE generator capacity by imposing a block diagonal structure on the generator that allows us to vary the added capacity.
-In the fourth part, we measure how much different models depend on the positional information in the image/video by shuffling the patches.
-A higher accuracy drop with randomly shuffled patches means the model relies more on the positions of the patches during inference.
-4.2.2 3D Classification
-In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-mixed and absolute encoding similar to the previous section [arnab2021vivit, heo2024rotary].
-For the 3D experiments, we examine video classification performance in the UCF101 dataset [soomro2012ucf101].
-Again, we did not optimize any hyperparameters for the LieRE model and used the dataloader from [tong2022videomae].
+In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters and use the same default hyperparameters for all experiments (Appendix B).
+We evaluate two versions of LieRE, distinguished by the basis matrix tile sizes of 64 and 8, referred to as LieRE64 and LieRE8, respectively.
+Notably, a tile size of 2 corresponds to RoPE-Mixed.
+5.1 2D Classification
+For 2D data we evaluate performance on the CIFAR-100 and ImageNet-1k image classification task (Krizhevsky et al., 2009; Deng et al., 2009).
+We further analyse different backbone models sizes and data ablation effects.
+We examine accuracy both CIFAR-100 and ImageNet-1k and compare to absolute (Dosovitskiy et al., 2020) position encoding and recent related work, RoPE-Mixed (Heo et al., 2024), VisionLlama embeddings (Chu et al., 2024).
+For 2D image classification tasks, we demonstrate that the LieRE-based transformer achieves a relative performance improvement in accuracy of 10.0% over standard absolute position encoding (Touvron et al., 2022), 7.3% over RoPE adaptation in VisionLlama (Chu et al., 2024), and 2.2% over RoPE-Mixed (Heo et al., 2024) on CIFAR-100, with similar accuracy trends observed on ImageNet (Deng et al., 2009) (table 1).
+Table 1: 2D image and 3D video classification Top-1 Accuracy (95% confidence intervals) results.
+All models use 85.2M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Stein et al., 2019) equivalent to DeiT architecture, ∗∗ equivalent to Vivit (spatio-temporal) architecture.
+Method CIFAR-100 ImageNet-1k UCF101
+Abs. Pos. E.∗,∗∗ 63.9 (62.9-65.8) 66.1 (65.7-66.5) 40.9 (40.5-41.3)
+VisionLlama RoPE 65.5 (64.6-66.5) 65.4 (65.0-65.8) 45.0 (44.6-45.4)
+RoPE-Mixed 68.8 (67.9-69.7) 68.8 (68.4-69.2) 46.3 (45.9-46.7)
+LieRE8 70.3 (69.4-71.2) 69.6 (69.2-70.0) 47.0 (46.6-47.4)
+LieRE64 70.0 (69.1-70.9) 69.3 (68.9-69.7) 44.7 (44.3-45.1)
+We vary the capacity of the transformer backbone with corresponding to ViT-Tiny, ViT-B and ViT-L, as proposed in (Dosovitskiy et al., 2020).
+Additionally, we analyze the impact of incorporating LieRE on performance across different model sizes on the CIFAR-100 dataset, as shown in table 5.
+Our results demonstrate that LieRE8 slightly outperforms alternatives across all evaluated model sizes.
+Figure 5: Performance behavior on CIFAR-100 (2D Image Classification) over ViT-Tiny (22M), ViT-Base (85M), ViT-Large (302M) for LieRE Rope-Mixed and Absolute Encoding (Appendix, table 5).
+Our experiments reveal a pattern in how learnable relative position encodings perform under limited data conditions.
+Specifically, when we varied the available training data from 20% to 90% of the CIFAR-100 dataset, both LieRE and RoPE-Mixed demonstrated significantly better data efficiency compared to traditional transformer approaches for 2D image classification (figure 6).
+Figure 6: Performance comparison of different positional embedding methods across varying dataset sizes.
+The plot shows the peak accuracy achieved by LieRE8, LieRE64, RoPE-Mixed, VisionLlaMA, and Absolute PE when trained on different fractions of the CIFAR-100 dataset.
+Both LieRE variants and RoPE-Mixed consistently outperform other methods, with their advantage becoming particularly pronounced in data-scarce scenarios.
+5.2 Synthetic Spatial Classification
+Figure 7: The model is asked to identify the direction of the arrow pointed to by the base of the Y. Here said arrow points down.
+Recent observations indicate that even advanced models like ChatGPT-4 and Claude Sonnet 3.5 struggle with basic spatial reasoning tasks.
+To investigate whether these limitations stem from positional encoding mechanisms, we designed a synthetic image classification task (Shah et al., 2024).
+The task presents a $108\times 108$ pixel image containing a $9\times 9$ grid (81 cells).
+Within this grid, we randomly place eight arrows and six letters (A, B, C, D, E, and Y), ensuring that one arrow is placed in the direction of the base of the ”Y”.
+The objective is to identify the direction of this specific arrow.
+To introduce visual complexity, we include four spurious letters and seven additional arrows as distractors.
+Figure 7 illustrates an example of this task setup.
+We verified that none of ChatGPT 4o, Claude Sonnet 3.5 and Gemeni Pro 1.5 are able to solve this task.
+We train the models on 800,000 examples and observe that they generally converge after the first 400,000.
+The only exception to this is absolute position encodings, where we have variants trained on 800,000 and
+Table 2: Comparing different positional encodings on synthetic task on base model (85M) equivalent to DeiT architecture across different resolutions, $108\times 108$, $168\times 168$, $276\times 276$
+Method 108 168 276
+Abs. Pos. E. 45.1 (43.1-47.7) 41.0 (39.0-43.2) 40.1 (38.2-42.0)
+Abs. Pos. E. (2M examples) 47.2 (45.2-49.1) - -
+VisionLlama RoPE 46.4 (44.0-47.9) 48.6 (46.0-49.9) 48.6 (46.2-50.1)
+RoPE-Mixed 100 (99.5-100) 98.6 (97.6-98.7) 88.6 (87.4-89.9)
+LieRE8 99.5 (99.2-99.8) 99.7 (99.4-99.9) 99.7 (99.5-99.9)
+LieRE64 100 (99.5-100) 100 (99.6-100) 100 (99.7-100)
+Table 2 outlines the performance of different positional encoding methods on a synthetic task using a base model (85M parameters) equivalent to the DeiT architecture.
+We evaluate the models across three different input resolutions ($108\times 108$, $168\times 168$, and $276\times 276$ pixels), revealing differences in scalability and effectiveness.
+While absolute positional encoding shows degraded performance as resolution increases (from 45.1% to 40.1%) and RoPE-Mixed demonstrates strong but degrading performance at higher resolutions (from 100% to 88.6%), both LieRE variants maintain near-perfect accuracy across all resolution scales, with LieRE64 achieving 100% accuracy consistently, suggesting superior resolution invariance of our Lie group-based approach.
+Qualitative analysis shows that absolute and visionllama position encoding attend less clearly to the ”Y” token than Rope-Mixed and LieRE.
+Please refer to the appendix for attention map examples, Figure 15 and Figure 15.
+5.3 3D Classification
+In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-Mixed and absolute encoding similar to the previous section (Arnab et al., 2021; Heo et al., 2024).
+For the 3D experiments, we examine video classification performance in the UCF101 dataset (Soomro et al., 2012).
+Again, we did not optimize any hyperparameters for the LieRE model and used the dataloader from (Tong et al., 2022).
 The full set of hyperparameters may be found in appendix B.
-5 Results
-5.1 Accuracy
-For 2D images, we show that the LieRE-based transformer outperforms the DeiT by a relative performance improvement of 9.4% [touvron2022deit], RoPE adaptation in VisionLlama [chu2024visionllama] by 6.7% and RoPE-Mixed by 9.7% on CIFAR-100 [heo2024rotary].
-For 3D inputs (video), we observe an relative accuracy improvement of LieRE-based transformer over the absolute position embedding of up to 25.5 % and RoPE-inspired position encodings at least to 6.1% in accuracy.
-Table 1: 2D image and 3D video classification Top-1 Accuracy (95% confidence intervals) results.
-All models use 85.1M parameters for 2D tasks and 88.7M parameters for 3D task [krizhevsky2009learning, deng2009imagenet, soomro2012ucf101, rsna2019data] equivalent to DeiT, ∗∗ equivalent to Vivit (spatio-temporal).
-Method CIFAR-100 ImageNet-1k UCF101
-Abs. Pos. E.∗,∗∗ 63.9 (62.9-65.8) 66.1 (65.7-66.5) 42.7 (42.3-43.1)
-VisionLlama RoPE 65.5 (64.6-66.5) 65.4 (65.0-65.8) 50.5 (50.1-50.9)
-RoPE-Mixed 63.7 (61.8-65.6) 65.9 (65.5-66.3) 45.7 (45.2-46.0)
-LieRE-Commute 66.9 (66.0 - 67.9) 68.4 (67.9-68.9) 52.1 (51.7-52.5)
-LieRE 69.9 (68.9-70.8) 69.5 (69.1-69.9) 53.6 (53.2-54.0)
-5.2 Data efficiency
-We observe that transformers based on LieRE exhibit greater data efficiency compared to leading transformer architectures for 2D images on the CIFAR-100 dataset.
-For instance, when using 40% of the dataset, the LieRE-based transformer outperforms all comparable methods by 4.5 percentage points.
-Additionally, there is a mild trend indicating that the performance gap widens as the amount of data decreases (Figure 3(b)).
-(a) Performance scales with generator dimension the number of LieRE parameters, compared to LieRE with one skew-symmetric matrix learned for the entire model (stem).
-(b) Data ablation on for different position embeddings on CIFAR-100.
-5.3 Model Scaling
-We examine three axes of capacity scaling: transformer backbone parameters, LieRE generator parameters and learning separate LieRE generators across heads and layers.
-Here we find a surprising result that RoPE-mixed can be substantially improved by sharing parameters across attention heads and layers and not learning them separately.
-5.3.1 Transformer backbone capacity
-We also evaluate how the inclusion of LieRE affects performance across model sizes on the CIFAR-100 in Figure 4(a).
-We observe that LieRE retains a statistically significant lead in performance across all three model sizes.
-Refer to caption (a) Performance behavior on CIFAR-100 (2D Image Classification) over ViT-Tiny (22M), ViT-Base (85M), ViT-Large (302M) for LieRE Rope-Mixed and Absolute Encoding (Appendix, table 5).
-Refer to caption (b) The LieRE spatial encoding allows the model to match the performance of absolute position encodings with substantially less training time.
-5.3.2 Generator parameters
-LieRE adds a small amount of capacity to the model ($4032$ parameters for the ViT-B backbone we use for most experiments), leading to the natural question of how helpful the marginal capacity is.
-We vary capacity by enforcing a block diagonal structure on the generator.
-varying the block size allows us to approach "full" LieRE.
-Recall that using $2\times 2$ blocks recovers exact commutativity and is referred to as LieRE-commute.
-This configuration substantially outperforms RoPE-Mixed, despite differing only in that it uses the same parameters across heads and layers and RoPE-Mixed learns separate parameters.
-In figure 3(a) we evaluate accuracy versus block dimension. We observe the steepest increase at the very start, though accuracy continues to grow as the generator dimension increases.
-5.3.3 Sharing LieRE Parameters Across Heads and Layers
-We also examine whether learning separate positional encoding per attention head and per layer improves performance on CIFAR-100 in table 2.
-We conduct this experiment for both RoPE-Mixed and LieRE and observe that in both cases it is best to learn one set of parameters across heads and layers.
-This is the mechanism by which we are able to show larger improvements from adopting RoPE-Mixed than the original paper [heo2024rotary].
-We name this transformed form of RoPE-Mixed LieRE-Commute, since it is equivalent to LieRE with generator dimension 2 (see Figure LABEL:fig:generator_dataset).
-Table 2: Layer, Head and Stem Learnable Rotary Embeddings on CIFAR-100, ViT-B, equivalent to LieRE-Commute
-FLOP Stem Heads Layers Rope-Mixed LieRE Reference
-5.684G 63.0 64.5 [heo2024rotary]
-5.684G 63.6 65.1
-5.613G 68.3 69.9
-5.613G 66.9 69.9 (ours)
-5.4 Compute efficiency
-Training transformers can necessitate substantial computational resources, which can hinder equitable access to research and development of machine learning methods.
-We demonstrate that the LieRE-based transformer requires 3.5 times less training epochs on Cifar100 to achieve comparable performance to the Absolute Position Embedding baseline (as used in DeiT III [touvron2022deit]).
-This represents the largest reduction in training time compared to recent works such as VisionLlama and RoPE-Mixed [chu2024visionllama, heo2024rotary].
-We note that this is the first training efficiency comparison of these recent methods.
-Figure 4(b) shows the amount of allowable compute reduction to achieve the same accuracy achieved by absolute position encodings (DeiT baseline) after 200 epochs.
-LieRE demonstrates the largest win, allowing a 3.5X reduction in training compute to achieve the same accuracy.
-5.5 FLOP analysis
-We find that since all methods we examine introduce a computational cost that is at most linear in the number of tokens, and runtime is dominated by the quadratic attention component, there is no substantial difference in computational efficiency between the methods.
-We list inference FLOP of the various methods in table 3.
-Table 3: FLOP analysis with percentage increase compared to absolute position encodings
-Position Enc. ViT-Tiny (22M) ViT-Base (85M) ViT-Large (302M)
-Abs. Pos. E. 0.963G 5.607G 19.856G
-VisionLlaMA RoPE 0.963G (+0.001%) 5.607G (+0.002%) 19.856G (+0.000%)
-RoPE-Mixed 0.963G (+0.001%) 5.609G (+0.036%) 19.863G (+0.035%)
-LieRE-Commute (ours) 0.963G (+0.000%) 5.608G (+0.018%) 19.858G (+0.005%)
-LieRE (ours) 0.970G (+0.727%) 5.613G (+0.107%) 19.869G (+0.065%)
+We observe a relative accuracy improvement of the LieRE-based transformer of up to 15.1% compared to absolute position embeddings and at least 1.5% compared to RoPE-inspired position encodings (table 1).
+5.4 LieRE parameter
+LieRE adds a small amount of capacity to the model ($580k$ parameters 0.68% for the ViT-B backbone we use for most experiments), leading to the natural question of how helpful the marginal capacity is.
+To answer this question, we vary the density of the learned basis and the learnability across heads and layers.
+We vary capacity by enforcing a block diagonal structure on the skew-symmetric basis.
+Varying the block size allows us to approach LieRE64.
+Recall that using $2\times 2$ blocks recovers exact commutativity and is identical to RoPE-Mixed.
+As RoPE-Mixed is sensitive to the initialization of the weights, we initialize all generator weights mirroring the original presentation of RoPE-Mixed (Heo et al., 2024).
+In figure 8 we evaluate accuracy versus block dimension (Appendix Table 8).
+Figure 8: Performance varies with skew-symmetric basis learnable dimensions, referred to as LieRE parameters.
+This is equivalent to increasing the tile size in the skew-symmetric basis ($2\times 2,4\times 4,8\times 8,16\times 16,32\times 32,48\times 48$).
+For both 2D (CIFAR-100) and 3D (UCF101) LieRE with tile size $8\times 8$ performs superior.
+Table 3: Accuracy with parameter sharing over heads and layers for ViT-B sized models on CIFAR-100.
+FLOP All Shared Shared Across Layers Shared Across Heads RoPE-Mixed LieRE64 LieRE8
+5.684G 68.8 70.0 70.3
+5.684G 68.7 69.5 69.8
+5.613G 69.5 69.7 69.7
+5.613G 68.3 69.4 69.5
+Table 3 presents the impact of sharing LieRE parameters across attention heads and layers on CIFAR-100 classification performance for RoPE-Mixed, LieRE64 and LieRE8 .
+We evaluate whether learning separate positional encodings for each attention head and layer provides performance benefits.
+We observe that learning across heads and layers yields superior performance followed by learning across layers.
 5.5 Patch shuffling
@@
 We evaluate models using the decline in accuracy when evaluating on shuffled patches.
-We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE models are most capable at using positional information.
+We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE models are more capable at using positional information.
 The complete results are displayed for CIFAR-100 and table for UCF101 (table 4).
@@
 […]
-Abs. Pos. E. 63.9 19.6 69.3 42.7 42.5 0.0
-VisionLlama RoPE 65.8 29.7 54.8 50.5 36.2 28.5
-RoPE-Mixed 63.6 38.7 39.0 45.7 37.9 17.1
-LieRE-Commute 66.9 16.5 75.4 52.1 37.2 28.6
-LieRE 69.9 15.9 77.1 53.6 24.7 53.9
-6 Limitations
-While LieRE shows promising results across multiple modalities and input dimensionalities, there are a few limitations worth noting.
-Our method is specifically designed to modify the inner product, making it compatible with most attention schemes, including original attention and linear attention.
-However, this specificity may limit its applicability to other architectures, like convolutional neural networks, that do not rely on inner product-based attention mechanisms.
-Future work could explore adaptations of our method to a wider range of architectures.Furthermore, the current formulation of our method is designed to encode vector positions in $\mathbb{R}^{d}$.
-While this is sufficient for many applications, it may not be directly applicable to tasks that require encoding poses in $SE(3)$ like robotics.
-Further research may be necessary to adapt our method to effectively handle such representational requirements.
-Despite these limitations, we believe that our work provides much-needed insight into how to improve model performance and reduce training costs by encoding relative position information across various modalities and input dimensionalities.
-7 Broader Impacts
-LieRE, our proposed method for encoding positional information in attention mechanisms, has demonstrated improvements in 2D image and video classification tasks.
-We are particularly excited about how LieRE can expand the applicability transformers to $n$ dimensional inputs.
-This may lay down the tracks to apply that same relative position encoding within and across modalities and settings.
+Abs. Pos. E. 63.9 19.6 69.3 40.9 39.5 0.0
+VisionLlama RoPE 65.5 29.7 54.8 45.0 37.0 17.7
+RoPE-Mixed 68.8 17.1 75.1 46.3 28.2 39.1
+LieRE8 70.3 12.3 82.5 47.0 27.8 40.9
+LieRE64 70.0 10.8 84.6 44.7 28.0 37.4
+Figure 9: Validation accuracy comparison between LieRE 64 and RoPE-Mixed positional embeddings across different image resolutions on ImageNet.
+Left: Models trained at $224\times 224$ resolution.
+Right: Models pre-trained at $224\times 224$ and fine-tuned at $256\times 256$ resolution.
+Both approaches show similar performance trends up to $320\times 320$, after which LieRE 64 demonstrates significantly better accuracy retention at higher resolutions, particularly in the fine-tuned scenario.
+5.6 Multi-resolution Classification
+In this section we compare the ability of methods to generalize to image resolutions not seen during training.
+We evaluate two training recipes inspired by (Heo et al., 2024).
+The first recipe matches the rest of the paper and consists of training the models on images of size $224\times 224$ for 200 epochs.
+The second adds an additional fine-tuning step at size $256\times 256$ for 30 epochs.
+The full details can be found in appendix B.
+We evaluate the accuracy on the ImageNet validation set with varying inference resolutions.
+Specifically, we scale the input images to resolutions of $196\times 196$, $256\times 256$, $320\times 320$, $384\times 384$, and $448\times 448$ pixels per dimension, and present the resulting accuracies in figure 9.
+For position assignment, we adopt a sequential approach where token positions are scaled proportionally to the image dimensions.
+For example, doubling the length of an image in each dimension doubles the range of positional indices.
+This method outperforms rescaling positions to a fixed range, as demonstrated by superior results for both RoPE-Mixed and LieRE across the evaluated training recipes.
 6 Conclusion
@@
 By requiring no changes to the tokenizer other than outputting positions and no other code changes, LieRE may provide a unified and efficient approach for transformers to process and learn from various data modalities within a single architecture.
-[…]
+7 Limitations
+While LieRE shows promising results across multiple modalities and input dimensionalities, there are a few limitations worth noting.
+Our method is specifically designed to modify the inner product, making it compatible with most attention schemes, including original attention and linear attention.
+However, this specificity may limit its applicability to other architectures, like convolutional neural networks, that do not rely on inner product-based attention mechanisms.
+Future work could explore adaptations of our method to a wider range of architectures.
+Furthermore, the current formulation of our method is designed to encode vector positions in $\mathbb{R}^{d}$.
+While this is sufficient for many applications, it may not be directly applicable to tasks that require encoding poses in $SE(3)$ like robotics.
+Further research may be necessary to adapt our method to effectively handle such representational requirements.
+Despite these limitations, we believe that our work provides much-needed insight into how to improve model performance and reduce training costs by encoding relative position information across various modalities and input dimensionalities.
+8 Broader Impacts
+LieRE, our proposed method for encoding positional information in attention mechanisms, has demonstrated improvements in 2D image and video classification tasks.
+We are particularly excited about how LieRE can expand the applicability transformers to $n$ dimensional inputs.
+This may lay down the tracks to apply that same relative position encoding within and across modalities and settings.
 Appendix A Appendix
@@
 Further experiments revealed substantial performance improvement with mean pooling and LieRE.
-We use the pytorch lightning framework for all experiments [falcon2019pytorch].
+We use the pytorch lightning framework for all experiments (Falcon, 2019).
 B.1 2D Image Classification
 The CIFAR experiments where trained on 8xL4 GPUs with 24GB of VRAM each and all took under 30 minutes to complete.
-The generator scaling experiment was conducted using RTX6000 GPUs.
+The basis capacity scaling experiment was conducted using RTX6000 GPUs.
 The ImageNet experiments were trained on 8xL40 GPUs and all took less than 2 days and 5 hours of runtime including time lost due to preemption and resource sharing.
@@
 We use a patch size of $4\times 4$ on the original $32\times 32$ image for CIFAR-100 and a patch size of $16\times 16$ on the randomly cropped and resized $224\times 224$ image.
-All vision experiments used RandAugment [cubuk2020randaugment].
+All vision experiments used RandAugment (Cubuk et al., 2020).
 We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$.
@@
 The 3D classification experiments were conducted on either $1008\times A$ 100 40GB GPUs or $1004\times A$ 100 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size.
-Similar to 2D classification, we use an initial learning rate of $41E-$ 4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $2\times 16\times 16$ on the randomly cropped and resized $32\times 224\times 224$ video/image.
+Similar to 2D classification, we use an initial learning rate of $41E-$ 4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $2\times 16\times 16$ on the randomly cropped and resized $8\times 224\times 224$ video/image.
 We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$.
-B.3 Scaling Model Size
-Table 5: Comparison of Position Encoding Methods for Different ViT Models, Accuracy (bootstrapped 95%CI)
-Position Encoding ViT-Tiny ViT-Base ViT-Large
+B.3 Multi-resolution Classification
+The second training recipe consists of 30 epochs with an initial learning rate of 1E-5 with a cosine decay.
+This mirrors the DEIT III training reciple that first pretrains at a lower resolution and finetunes at a higher resolution.
+B.4 CIFAR-100 Performance Across Model Scales
+We also evaluate how the inclusion of LieRE affects performance across model sizes on the CIFAR-100 in table 5.
+We observe that LieRE retains a statistically significant lead in performance across all three model sizes.
+B.4.1 Basis parameters
+Table 5: Comparison of Position Encoding Methods for Different ViT Models Sizes on CIFAR-100, Accuracy (bootstrapped 95%CI)
+Position Enc. ViT-Tiny (22M) ViT-Base (86M) ViT-Large (302M)
 Abs. Pos. E. 57.2 (56.2-58.1) 63.9 (62.9-65.8) 60.5 (59.5-61.4)
 VisionLlaMA RoPE 58.2 (57.2-59.2) 65.5 (64.6 -66.5) 62.3 (60.4-64.2)
-RoPE-Mixed 59.6 (58.7 - 60.6) 63.7 (61.8 - 65.6) 65.0 (64.0 - 65.9)
-LieRE-Commute 63.2 (62.2 - 64.1) 66.9 (66.0 - 67.9) 65.1 (64.1 - 66.0)
-LieRE 64.7 (63.8 - 65.7) 69.9 (68.9 - 70.8) 69.2 (68.3 - 70.1)
-B.4 Python implementation of skew-symmetric matrix
-1generator_raw_params = nn.Parameter(
+RoPE-Mixed 65.4 (64.5-66.4) 68.8 (67.9-69.7) 68.8 (67.9-69.7)
+LieRE8 65.6 (64.7-66.6) 70.3 (69.4-71.2) 69.9 (68.9-70.8)
+LieRE64 65.3 (64.4-66.3) 70.0 (69.1-69.7) 68.9 (68.0-69.8)
+B.5 Python implementation of LieRE rotation matrix computation
+1basis_raw_params = nn.Parameter(
 2 torch.rand(
@@
 9upper_triangle = (
-10 torch.triu(generator_raw_params, diagonal=1)
+10 torch.triu(basis_raw_params, diagonal=1)
 11)
@@
 15)
-16generator_pos = torch.sum(in_basis_positions, dim=-3)
-17rotation = torch.matrix_exp(generator_pos.to(dtype=torch.float32)).to(dtype=positions.dtype)
+16rotation_log = torch.sum(in_basis_positions, dim=-3)
+17rotation = torch.matrix_exp(rotation_log.to(dtype=torch.float32)).to(dtype=positions.dtype)
+B.6 Compute efficiency
+We demonstrate that LieRE-based transformer achieves comparable performance to the Absolute Position Embedding baseline (DeiT III (Touvron et al., 2022)) on CIFAR-100 with fewer training epochs.
+This represents a notable advancement over recent methods such as VisionLlama and RoPE-Mixed (Chu et al., 2024; Heo et al., 2024).
+Figure 10 illustrates that LieRE enables a 3.9$\times$ reduction in training compute while maintaining the accuracy achieved by absolute position encodings after 200 epochs.
+Figure 10: The LieRE spatial encoding allows the model to match the performance of absolute position encodings with substantially less training time.
+B.7 FLOPS Comparison of methods
+We find that since all methods we examine introduce a computational cost that is at most linear in the number of tokens, and runtime is dominated by the quadratic attention component, there is no substantial difference in computational efficiency between the methods.
+We list inference FLOP of the various methods in table 6.
+Table 6: FLOP analysis with percentage increase compared to absolute position encodings
+Position Enc. ViT-Tiny (22M) ViT-Base (85M) ViT-Large (302M)
+Abs. Pos. E. 0.963G 5.607G 19.856G
+VisionLlaMA RoPE 0.963G (+0.001%) 5.607G (+0.002%) 19.856G (+0.000%)
+RoPE-Mixed 0.964G (+0.104%) 5.609G (+0.036%) 19.863G (+0.035%)
+LieRE8 0.968G (+0.519%) 5.617G (+0.178%) 19.882G (+0.065%)
+LieRE64 0.970G (+0.727%) 5.684G (+1.375%) 20.061G (+1.033%)
+B.8 Validation Losses
+Table 7: 2D image and 3D video classification Top-1 Validation loss (95% confidence intervals) results.
+All models use 85.1M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Stein et al., 2019) equivalent to DeiT, ∗∗ equivalent to Vivit (spatio-temporal).
+Method CIFAR-100 ImageNet-1k UCF101
+Abs. Pos. E.∗,∗∗ 1.56 (1.47-1.56) 1.84 (1.81-1.86) 2.94 (2.92-2.96)
+VisionLlama RoPE 1.56 (1.51-1.61) 1.98 (1.94-2.01) 2.66 (2.63-2.69)
+RoPE-Mixed 1.38 (1.33-1.43) 1.72 (1.68-1.74) 2.52 (2.49-2.54)
+LieRE8 1.36 (1.31-1.41) 1.73 (1.70-1.76) 2.47 (2.44-2.49)
+LieRE64 1.37 (1.33-1.42) 1.73 (1.70-1.76) 2.64 (2.62-2.67)
+B.9 Basis parameters
+Table 8: Accuracy Results for Different $\mathrm{LieRE}_{\Theta}$ Parameters, relative to the model size of 85.2M and 88.7M
+Dataset $\mathrm{LieRE}_{\Theta}$ Parameter Absolute $\mathrm{LieRE}_{\Theta}$ Parameter Relative Tile Size Accuracy (%) CI (95%)
+CIFAR100 9216 0.01 % 2 68.84 (67.93-69.75)
+CIFAR100 27648 0.03 % 4 69.28 (68.38-70.18)
+CIFAR100 64512 0.08 % 8 70.32 (69.42-71.22)
+CIFAR100 138240 0.16 % 16 69.85 (68.95-70.75)
+CIFAR100 285696 0.34 % 32 69.65 (68.75-70.55)
+CIFAR100 580608 0.68 % 64 69.99 (69.09-70.89)
+UCF101 9216 0.01 % 2 46.30 (45.89-46.71)
+UCF101 27648 0.03 % 4 45.67 (45.26-46.08)
+UCF101 64512 0.08 % 8 47.03 (46.62-47.44)
+UCF101 138240 0.16 % 16 46.86 (46.45-47.27)
+UCF101 285696 0.32 % 32 46.42 (46.01-46.83)
+UCF101 580608 0.66 % 64 44.68 (44.27-45.09)
+B.10 Validation Accuracy Curves
+Figure 11: CIFAR-100
+B.11 Training Loss Curves
+Figure 12: CIFAR-100
+B.12 Attention Maps
+We show the attention scores to the CLS token, averaged across heads for every layer.
+Each grid is rescaled so that the minimum element has value zero and maximum value of one.
+Red indicates the maximal score, and blue indicates the minimal value.
+Figure 13: Attention Maps (normalized), $108\times 108$: Rope-Mixed, LieRE, Absolute and Visionllama positional encoding (x-axis), Layers 1-12 (y-axis).
+While Rope-Mixed and LieRE learn to look at the ”Y”, Absolute and Visionllama less and concentrate on the arrows.
+Figure 14: Attention Maps (normalized), $276\times 276$: Rope-Mixed, LieRE, Absolute and Visionllama positional encoding (x-axis), Layers 1-12 (y-axis).
+While LieRE still learns to look at the ”Y”, Absolute and Visionllama do less so.
+Figure 15: Comparison of attention maps across different resolutions and position encoding methods.

### v4

@@
 Abstract
-Transformer architectures rely on position encodings to capture token dependencies.
+Transformer architectures depend on explicit position encodings to capture token positional information.
 Rotary Position Encoding (RoPE) has emerged as a popular choice in language models due to its efficient encoding of relative position information through key-query rotations.
 However, RoPE faces significant limitations beyond language processing: it is constrained to one-dimensional sequence data and, even with learnable phases, offers limited representational capacity.
-We address these challenges with Lie Relative Encodings (LieRE), which replaces RoPE’s block-2D rotation matrix with a learned, dense, high-dimensional rotation matrix of variable density.
-Through extensive evaluation on three image datasets across 2D and 3D classification tasks, LieRE achieves 2% relative improvement over state-of-the-art baselines on 2D tasks and 1.5% on 3D tasks, while demonstrating superior generalization to higher resolutions.
-Our implementation is computationally efficient, with results reproducible on 4 A100 GPUs in 30 minutes on CIFAR100, and we release our code to facilitate further research.
+We address these challenges with Lie Relative Encodings (LieRE), which generalizes RoPE to high-dimensional rotation matrices by leveraging their Lie group structure.
+Through extensive evaluation on three image datasets across 2D and 3D classification tasks, LieRE achieves 1.5% improvement over state-of-the-art baselines on 2D tasks and 1% on 3D tasks, while demonstrating superior generalization to higher resolutions.
+Our implementation is computationally efficient, with results reproducible on 4 A100 GPUs in 30 minutes on CIFAR100.
+Our code is available at https://github.com/StanfordMIMI/LieRE.
 Machine Learning, ICML
@@
 The attention mechanism, particularly within transformer architectures, has revolutionized machine learning across diverse domains.
-However, attention is inherently permutation—invariant—it cannot leverage the sequential order of its inputs as information.
-This fundamental limitation necessitates additional mechanisms to encode positional information, enabling models to capture sequential and spatial dependencies crucial for tasks ranging from natural language processing to image understanding (Vaswani et al., 2017).
+However, attention is inherently permutation invariant—it cannot leverage the sequential order of its inputs as information directly.
+This fundamental limitation necessitates additional mechanisms to encode positional information, enabling models to capture sequential and spatial dependencies crucial for tasks ranging from natural language to image understanding (Vaswani et al., 2017).
 This challenge has sparked extensive research into positional encodings, which inject order information into the otherwise order-agnostic attention mechanism.
@@
 This translation-invariant property makes RoPE particularly efficient at capturing position-dependent patterns in text, which has led to its adoption in popular open-source language models such as LLaMA and Mixtral.
-Figure 1: Overview of LieRE, where $A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x\;y\;z])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the first patch in the top left corner of the input with the position $P_{i}=(x,y,z)\in\mathbb{R}^{N}$.
+Figure 1: Overview of how LieRE encodes spatial information, where $N$ denotes the number of input dimensions, i.e.
+$N=3$ for the 3D image.
+$A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x_{i}\;y_{i}\;z_{i}])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the $i$th patch in the flattened input, where $j$ corresponds to a different patch.
 $d$ is the head dimension.
-The upper triangle contains learnable $\textrm{LieRE}_{\theta}$ parameter, while the lower triangle contains their negatives, -$\textrm{LieRE}_{\theta}$, reflecting the skew-symmetric nature of the matrix.
 Despite RoPE’s success in sequential tasks (Touvron et al., 2023; Chowdhery et al., 2023), it faces several fundamental limitations.
@@
 To assess the impact of LieRE and other positional encodings on ViT performance, we evaluate several encoding schemes across diverse tasks, including 2D and 3D image classification.
-Additionally, we investigate a fundamental spatial reasoning task where models must identify arrow directions.
+Additionally, we investigate a fundamental spatial reasoning task where models must identify where an arrow points.
 Despite their sophistication, contemporary multimodal LLMs struggle with this seemingly simple task.
@@
 2.1 Position Encodings
-We organize our literature review into three broad classes of positional encodings: 1) absolute, 2) relative, and 3) contextual.
-Absolute position in this context refers to a position with respect to a consistent reference, usually the start of a tex or top left corner of an image.
+We split the review of positional encodings into: a) absolute, b) relative, and c) contextual.
+Absolute position in this context refers to a position with respect to a consistent reference, usually the start of a text or top left corner of an image.
 Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the input.
 Methods such as sinusoidal and learned absolute encodings directly add vectors to the input token embedding (Vaswani et al., 2017; Devlin et al., 2019; Dosovitskiy et al., 2020).
-Relative position encodings in contrast encode the relative positions of two tokens.
+Relative position encodings, in contrast, encode the relative positions of two tokens.
 One strategy is to learn an set of embeddings for position deltas which can be incorporated into the attention mechanism (Shaw et al., 2018b; Liu et al., 2021, 2022).
@@
 The primary difference is that RoPE-Mixed has a learnable component, whereas VisionLlama does not.
+2.3 Efficient Scaling beyond Sequence Data
+There is also an extensive line of work improving the performance of transformers for modalities with dimensionality greater than one.
+Axial Attention (Ho et al., 2019) reduces computational complexity by applying attention along specific axes (e.g., rows and columns in images), enabling transformers to scale efficiently with high-dimensional data.
+Perceiver utilizes latent tokens to compress high-dimensional inputs into a smaller set, improving scalability as input size and dimensionality grow.
+These methods address the inefficiencies of traditional transformers when applied to high-dimensional data (Jaegle et al., 2021).
+Additionally, techniques like Swin and Vmamba optimize compute for visual data through structuring how information flows through the network (Liu et al., 2021, 2024).
+Swin Transformer introduces a hierarchical approach with shifted windows, limiting attention to local regions to reduce complexity while capturing global context.
+Vmamba, on the other hand, proposes a visual state space model that represents images as a collection of spatial states, allowing attention to be applied efficiently across large-scale visual inputs by exploiting spatial locality and reducing redundant computation.
+2.4 Equivariant Networks
+A related branch of work encoding problem structure focuses on equivariance.
+We say that a model $T$ is equivariant with respect to if $T(f(x))=g(T(x))g$ (Tai et al., 2019).
+Where with relative position encoding we often want to be able to encode translation invariance, equivariance provides a more general framework.
+Equivariance has been applied to improve performance on problems with a wide array of structures, ranging from rotation-invariance (Esteves et al., 2017, 2018; Worrall et al., 2017), 3D reference frame-invariance (Liao & Smidt, 2022; Fuchs et al., 2020) and many others.
+The subset of these works that focus on generating equivariant token embeddings for transformers can be combined directly with LieRE or another rotation-based position encoding.
+2.5 Lie Groups in Machine Learning
+Lie groups have also had extensive use in machine learning.
+The range of works is diverse, ranging from algebraic signal processing (Kumar et al., 2024), automated discovery of symmetries (Forestano et al., 2023) to state estimation (Falorsi et al., 2019).
+Furthermore (Gallier & Quaintance, 2020) provides a friendly introduction to differential geometry and lie groups that may be of interest to the reader.
 3 Background
@@
 $\exp(U-V)=\exp(-V+U)\approx\exp(V)^{-1}\exp(U)$ (1)
-Both RoPE (in the context of text) and RoPE-Mixed use block-diagonal rotation matrices with 2D rotations as blocks 4.
+Both RoPE (in the context of text) and RoPE-Mixed use block-diagonal rotation matrices with 2D rotations as blocks 3(b).
 These form a special Lie group that is commutative, allowing us to strengthen the statement in (1) to
@@
 Our work examines the tradeoff between using the stronger property in (2) or increased capacity and the weaker property (1).
-Figure 3: Pseudocode for the Lie Rotary Embedding (LieRE) attention mechanism. Unlike RoPE, LieRE allows for dense rotation matrices.
+(a) Lie Rotary Embedding (LieRE) attention mechanism.
+(b) Rotary Position Embedding (RoPE) attention mechanism.
+Figure 3: Comparison of the LieRE and RoPE-Mixed attention mechanisms.
 3.2 Attention Mechanism
@@
 4 Method
-LieRE is a simple modification to the attention mechanism that is presented in algorithm 3.
+LieRE is a simple modification to the attention mechanism that is presented in Algorithm 3(a).
 Recall that we assume that positions are $n$-dimensional vectors, a matrix $A$ is skew-symmetric if $A^{T}=-A$, and that the matrix exponential of a skew-symmetric matrix, call it $A$, is always a high dimensional rotation matrix.
@@
 Learning in the space of skew-symmetric matrices allows us to sidestep some of the difficulty that would come from learning on the manifold of rotation matrices (Figure 2).
-Figure 4: Pseudocode for the Rotary Position Embedding (RoPE) attention mechanism.
 LieRE uses the rotation matrix computed above to modify the keys and queries of the standard attention mechanism.
@@
 Note that the only difference between LieRE and RoPE-Mixed is that the latter constrains the rotations to be block-diagonal with block size two.
-We include the pseudocode for the LieRE attention in algorithm 3 in addition to the standard RoPE attention (algorithm 4).
+We include the pseudocode for the LieRE attention in Algorithm 3(a) in addition to the standard RoPE attention (Algorithm 3(b)).
 In practice, we compute the rotation matrices at the start of the forward pass.
 By default, the skew symmetric bases are learned separately for every layer and attention head except in the experimental section focused on sharing parameters across heads and layers.
+Table 1: 2D image and 3D video classification Top-1 Accuracy (95% confidence intervals) results.
+All models use 85.2M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Flanders et al., 2020) equivalent to DeiT architecture, ∗∗ equivalent to Vivit (spatio-temporal) architecture.
+Method CIFAR-100 ImageNet-1k UCF101
+Abs. Pos. E.∗,∗∗ 63.9 (62.9-65.8) 66.1 (65.7-66.5) 40.9 (40.5-41.3)
+VisionLlama RoPE 65.5 (64.6-66.5) 65.4 (65.0-65.8) 45.0 (44.6-45.4)
+RoPE-Mixed 68.8 (67.9-69.7) 68.8 (68.4-69.2) 46.3 (45.9-46.7)
+LieRE8 70.3 (69.4-71.2) 69.6 (69.2-70.0) 47.0 (46.6-47.4)
+LieRE64 70.0 (69.1-70.9) 69.3 (68.9-69.7) 44.7 (44.3-45.1)
+Table 2: P-values comparing RoPE-Mixed and LieRE8 across datasets.
+Dataset P-value (RoPE-Mixed vs. LieRE8)
+CIFAR-100 $1.3\times 10^{-5}$
+ImageNet-1k $6.3\times 10^{-3}$
+UCF101 $7.1\times 10^{-4}$
+$384\times 384$ (Resolution Invariance) $4.0\times 10^{-4}$
 Adjusting the skew-symmetric basis matrices’ block width allows us to incrementally adjust the capacity allocated towards position encoding.
@@
 We avoid using pre-trained weights in order to maximize the comparability of results between methods.
-In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters and use the same default hyperparameters for all experiments (Appendix B).
+In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters and use the same default hyperparameters for all experiments (Appendix B.1).
 We evaluate two versions of LieRE, distinguished by the basis matrix tile sizes of 64 and 8, referred to as LieRE64 and LieRE8, respectively.
@@
 We examine accuracy both CIFAR-100 and ImageNet-1k and compare to absolute (Dosovitskiy et al., 2020) position encoding and recent related work, RoPE-Mixed (Heo et al., 2024), VisionLlama embeddings (Chu et al., 2024).
-For 2D image classification tasks, we demonstrate that the LieRE-based transformer achieves a relative performance improvement in accuracy of 10.0% over standard absolute position encoding (Touvron et al., 2022), 7.3% over RoPE adaptation in VisionLlama (Chu et al., 2024), and 2.2% over RoPE-Mixed (Heo et al., 2024) on CIFAR-100, with similar accuracy trends observed on ImageNet (Deng et al., 2009) (table 1).
-Table 1: 2D image and 3D video classification Top-1 Accuracy (95% confidence intervals) results.
-All models use 85.2M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Stein et al., 2019) equivalent to DeiT architecture, ∗∗ equivalent to Vivit (spatio-temporal) architecture.
-Method CIFAR-100 ImageNet-1k UCF101
-Abs. Pos. E.∗,∗∗ 63.9 (62.9-65.8) 66.1 (65.7-66.5) 40.9 (40.5-41.3)
-VisionLlama RoPE 65.5 (64.6-66.5) 65.4 (65.0-65.8) 45.0 (44.6-45.4)
-RoPE-Mixed 68.8 (67.9-69.7) 68.8 (68.4-69.2) 46.3 (45.9-46.7)
-LieRE8 70.3 (69.4-71.2) 69.6 (69.2-70.0) 47.0 (46.6-47.4)
-LieRE64 70.0 (69.1-70.9) 69.3 (68.9-69.7) 44.7 (44.3-45.1)
+For 2D image classification tasks, we demonstrate that the LieRE-based transformer achieves a relative performance improvement in accuracy of 10.0% over standard absolute position encoding (Touvron et al., 2022), 7.3% over RoPE adaptation in VisionLlama (Chu et al., 2024), and 2.2% over RoPE-Mixed (Heo et al., 2024) on CIFAR-100, with similar accuracy trends observed on ImageNet (Deng et al., 2009) (Table 1, Table 2).
 We vary the capacity of the transformer backbone with corresponding to ViT-Tiny, ViT-B and ViT-L, as proposed in (Dosovitskiy et al., 2020).
-Additionally, we analyze the impact of incorporating LieRE on performance across different model sizes on the CIFAR-100 dataset, as shown in table 5.
+Additionally, we analyze the impact of incorporating LieRE on performance across different model sizes on the CIFAR-100 dataset, as shown in Table 6, Figure 8.
 Our results demonstrate that LieRE8 slightly outperforms alternatives across all evaluated model sizes.
-Figure 5: Performance behavior on CIFAR-100 (2D Image Classification) over ViT-Tiny (22M), ViT-Base (85M), ViT-Large (302M) for LieRE Rope-Mixed and Absolute Encoding (Appendix, table 5).
-Our experiments reveal a pattern in how learnable relative position encodings perform under limited data conditions.
-Specifically, when we varied the available training data from 20% to 90% of the CIFAR-100 dataset, both LieRE and RoPE-Mixed demonstrated significantly better data efficiency compared to traditional transformer approaches for 2D image classification (figure 6).
 Figure 4: Performance comparison of different positional embedding methods across varying dataset sizes.
@@
 Both LieRE variants and RoPE-Mixed consistently outperform other methods, with their advantage becoming particularly pronounced in data-scarce scenarios.
+Our experiments reveal a pattern in how learnable relative position encodings perform under limited data conditions.
+Specifically, when we varied the available training data from 20% to 90% of the CIFAR-100 dataset, both LieRE and RoPE-Mixed demonstrated significantly better data efficiency compared to traditional transformer approaches for 2D image classification (figure 4).
 5.2 Synthetic Spatial Classification
-Figure 7: The model is asked to identify the direction of the arrow pointed to by the base of the Y. Here said arrow points down.
+Figure 5: The model is asked to identify the direction of the arrow pointed to by the base of the Y. Here said arrow points down. This task requires understanding spatial relationships.
 Recent observations indicate that even advanced models like ChatGPT-4 and Claude Sonnet 3.5 struggle with basic spatial reasoning tasks.
@@
 We verified that none of ChatGPT 4o, Claude Sonnet 3.5 and Gemeni Pro 1.5 are able to solve this task.
-We train the models on 800,000 examples and observe that they generally converge after the first 400,000.
-The only exception to this is absolute position encodings, where we have variants trained on 800,000 and
+We train the models on 800,000 examples and observe that they generally converge after the first 400,000 examples.
+The only exception to this is absolute position encodings, where all variants were trained on 800,000 examples with no sign of increases in performance.
 Table 3: Comparing different positional encodings on synthetic task on base model (85M) equivalent to DeiT architecture across different resolutions, $108\times 108$, $168\times 168$, $276\times 276$
@@
 Again, we did not optimize any hyperparameters for the LieRE model and used the dataloader from (Tong et al., 2022).
-The full set of hyperparameters may be found in appendix B.
+The full set of hyperparameters may be found in appendix B.1.
 We observe a relative accuracy improvement of the LieRE-based transformer of up to 15.1% compared to absolute position embeddings and at least 1.5% compared to RoPE-inspired position encodings (table 1).
-5.4 LieRE parameter
+5.4 LieRE Capacity
 LieRE adds a small amount of capacity to the model ($580k$ parameters 0.68% for the ViT-B backbone we use for most experiments), leading to the natural question of how helpful the marginal capacity is.
-To answer this question, we vary the density of the learned basis and the learnability across heads and layers.
-We vary capacity by enforcing a block diagonal structure on the skew-symmetric basis.
-Varying the block size allows us to approach LieRE64.
+To answer this question, we vary the density of the learned basis and the learnability between heads and layers.
+We vary capacity by enforcing a block-diagonal structure on the skew-symmetric basis.
+Varying the block size allows us to approach a dense matrix (LieRE64 for our models).
 Recall that using $2\times 2$ blocks recovers exact commutativity and is identical to RoPE-Mixed.
 As RoPE-Mixed is sensitive to the initialization of the weights, we initialize all generator weights mirroring the original presentation of RoPE-Mixed (Heo et al., 2024).
+We found that RoPE-Mixed is more sensitive to the initialization method than LieRE as seen in appendix table 7.
 In figure 6 we evaluate accuracy versus block dimension (Appendix Table 10).
@@
 The second adds an additional fine-tuning step at size $256\times 256$ for 30 epochs.
-The full details can be found in appendix B.
+The full details can be found in appendix B.1.
 We evaluate the accuracy on the ImageNet validation set with varying inference resolutions.
@@
 Despite these limitations, we believe that our work provides much-needed insight into how to improve model performance and reduce training costs by encoding relative position information across various modalities and input dimensionalities.
-8 Broader Impacts
-LieRE, our proposed method for encoding positional information in attention mechanisms, has demonstrated improvements in 2D image and video classification tasks.
-We are particularly excited about how LieRE can expand the applicability transformers to $n$ dimensional inputs.
-This may lay down the tracks to apply that same relative position encoding within and across modalities and settings.
-Appendix A Appendix
-Appendix B Experiment Hyperparameters
+Impact statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Acknowledgements
+We would like to thank Maksim Maydanskiy for helping us understand Lie groups and Lie group representations.
+We would like to thank Aradhana Sinha for suggesting the shuffling experiment and providing early feedback on the paper.
+Sophie Ostmeier was supported by the German Research Foundation (DFG), Walter-Benjamin fellowship (ID: 517316550).
+This project was supported by Google Cloud and Azure credits.
+Appendix A Equivalence of RoPE and LieRE in one Dimension
+Though focused on higher dimensional inputs LieRE remains compatible with 1D tasks.
+It turns out that in the 1D setting, LieRE has identical representational capacity to RoPE.
+This is not the case for higher dimensional inputs
+Recall that in the 1D setting positions are scalars.
+The LieRE rotation is $R=\exp(tA)$ for a some learnable skew-symmetric matrix $A$.
+Recall that skew-symmetric matrices can be written in the form $S^{T}\Lambda S$ where and is orthogonal $\Lambda$ has the structure denoted below.
+$\Lambda=\begin{pmatrix}0&\lambda_{0}&&&&\\ -\lambda_{0}&0&&&&\\ &&0&\lambda_{1}&&\\ &&-\lambda_{1}&0&&\\ &&&&\ddots&\\ \end{pmatrix}$
+We can then use an identity of the matrix exponential to break down the LieRE rotation matrix.
+$R=\exp(tS^{T}\Lambda S)=S^{T}\exp(t\Lambda)S.$
+For two tokens in positions $t,t^{\prime}$ we denote the embeddings for a specific attention head as $x_{t},x_{t^{\prime}}$.
+If $K,Q$ denote the corresponding key and query linear transformation matrices we can write the attention inner product with LieRE explicitly.
+$x_{t}KR_{t}^{T}R_{t^{\prime}}Qx_{t^{\prime}}$ $=x_{t}KS^{T}\exp(t\Lambda)S^{T}S\exp(t^{\prime}\Lambda)SQx_{t^{\prime}}$
+$=x_{t}KS^{T}\exp(t\Lambda)^{T}\exp(t^{\prime}\Lambda)SQx_{t^{\prime}}$
+$=x_{t}KS^{T}\exp(t\Lambda)^{T}\exp(t^{\prime}\Lambda)SQx_{t^{\prime}}$
+We let $K^{\prime}=KS^{T}$ and $Q^{\prime}=SQ$, since these matrices are all learnable we can fold the $S$ matrix into parameters of the key and query linear layers for the given head, allowing us to simplify the above expression.
+$x_{t}K^{\prime}\exp(t\Lambda)^{T}\exp(t^{\prime}\Lambda)Q^{\prime}x_{t^{\prime}}$
+Now we use the fact that each block is skew symmetric. In the case of two dimensions,
+$\exp\left(\begin{pmatrix}0&\lambda\\ -\lambda&0\end{pmatrix}\right)=\begin{pmatrix}\cos(\lambda)&\sin(\lambda)\\ -\sin(\lambda)&\cos(\lambda)\end{pmatrix}$
+If we let $R_{\lambda}$ denote a block diagonal rotation matrices with 2D rotations of angles $\lambda_{0},\ldots,\lambda_{n}$, we can rewrite the above expression in a more familiar form.
+$x_{t}K^{\prime}R_{t}^{T}R_{t^{\prime}}Q^{\prime}x_{t^{\prime}}$
+This is exactly the formulation of the original RoPE position encoding.
+This also makes more clear how LieRE is different from RoPE-Mixed in the high dimensional setting.
+The above proof depends on the fact that we can decompose every rotation into a matrix of the form $S^{T}\Lambda S$ with $S$ not dependent on the position, allowing us to fold the orthogonal $S$ matrices into the key and query matrices.
+This decomposition with constant $S$ is guaranteed because the inputs to the matrix exponential differ by only a scalar factor.
+This is no longer true once we switch to more than a one-dimensional basis of skew symmetric matrices.
+Appendix B Experimental Details
+B.1 Experimental Hyperparameters
 The backbone for all experiments is configured as ViT-B, with 12 layers, a hidden dimension of 768, and an intermediate dimension of 3096.
@@
 We use the pytorch lightning framework for all experiments (Falcon, 2019).
-B.1 2D Image Classification
+B.2 2D Image Classification
 The CIFAR experiments where trained on 8xL4 GPUs with 24GB of VRAM each and all took under 30 minutes to complete.
@@
 The hyperparameters were tuned with RoPE-Mixed and selected before conducting the LieRE trainers as to ensure a fair comparison.
-B.2 3D Video Classifications
+B.3 3D Video Classifications
 The 3D classification experiments were conducted on either $1008\times A$ 100 40GB GPUs or $1004\times A$ 100 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size.
@@
 We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$.
-B.3 Multi-resolution Classification
+B.4 Multi-resolution Classification
 The second training recipe consists of 30 epochs with an initial learning rate of 1E-5 with a cosine decay.
-This mirrors the DEIT III training reciple that first pretrains at a lower resolution and finetunes at a higher resolution.
-B.4 CIFAR-100 Performance Across Model Scales
-We also evaluate how the inclusion of LieRE affects performance across model sizes on the CIFAR-100 in table 5.
-We observe that LieRE retains a statistically significant lead in performance across all three model sizes.
-B.4.1 Basis parameters
+This mirrors the DEIT III training recipe that first pretrains at a lower resolution and finetunes at a higher resolution.
+B.5 CIFAR-100 Performance Across Model Scales
+We further evaluate the impact of incorporating LieRE across different model sizes on CIFAR-100, as shown in Table 6.
+LieRE consistently outperforms the baseline with statistically significant gains across all three model scales.
+However, these results may be sensitive to dataset size, as all models are trained from scratch in this study.
+This is reflected in the performance drop observed with the ViT-Huge model.
+Figure 8: Performance behavior on CIFAR-100 (2D Image Classification) over ViT-Tiny (22M), ViT-Base (85M), ViT-Large (302M) for LieRE RoPE-Mixed and Absolute Encoding (Appendix, table 6).
 Table 6: Comparison of Position Encoding Methods for Different ViT Models Sizes on CIFAR-100, Accuracy (bootstrapped 95%CI)
@@
 LieRE64 65.3 (64.4-66.3) 70.0 (69.1-69.7) 68.9 (68.0-69.8)
-B.5 Python implementation of LieRE rotation matrix computation
-1basis_raw_params = nn.Parameter(
-2 torch.rand(
-3 input_dimensionality,
-4 head_dim,
-5 head_dim,
-6 ) * 2 * math.pi # optional, inspired from RoPE-Mixed paper
-7)
-8
-9upper_triangle = (
-10 torch.triu(basis_raw_params, diagonal=1)
-11)
-12skew_bases = upper_triangle - torch.transpose(upper_triangle, -1, -2)
-13in_basis_positions = (
-14 positions.reshape(list(positions.shape) + [1] * 2) * skew_bases
-15)
-16rotation_log = torch.sum(in_basis_positions, dim=-3)
-17rotation = torch.matrix_exp(rotation_log.to(dtype=torch.float32)).to(dtype=positions.dtype)
-B.6 Compute efficiency
+B.6 Initialization Sensitivity
+In order to understand the sensitivity of the methods to the initialization, we explore several scaling factor for the weights initialization. The results are presented in table 7.
+Table 7: Paired Z-test between $2\pi$ and $1$ initialization for LieRE8 and RoPE-Mixed.
+Metric LieRE8 ($2\pi$ vs 1 init) RoPE-Mixed ($2\pi$ vs 1 init)
+Z-statistic -1.81 -3.85
+P-value 0.070 0.00012
+Difference between means -0.0118 -0.0255
+95% Confidence Interval [-0.0246, 0.0010] [-0.0385, -0.0125]
+Appendix C Python implementation of LieRE rotation matrix computation
+basis_raw_params = nn.Parameter( torch.rand( input_dimensionality, head_dim, head_dim, ) * 2 * math.pi # optional, inspired from RoPE-Mixed paper ) upper_triangle = ( torch.triu(basis_raw_params, diagonal=1) ) skew_bases = upper_triangle - torch.transpose(upper_triangle, -1, -2) in_basis_positions = ( positions.reshape(list(positions.shape) + [1] * 2) * skew_bases ) rotation_log = torch.sum(in_basis_positions, dim=-3) rotation = torch.matrix_exp(rotation_log.to(dtype=torch.float32)) rotation = rotation.to(dtype=positions.dtype)
+Appendix D Compute efficiency
 We demonstrate that LieRE-based transformer achieves comparable performance to the Absolute Position Embedding baseline (DeiT III (Touvron et al., 2022)) on CIFAR-100 with fewer training epochs.
@@
 Figure 9: The LieRE spatial encoding allows the model to match the performance of absolute position encodings with substantially less training time.
-B.7 FLOPS Comparison of methods
+D.1 FLOPS Comparison of methods
 We find that since all methods we examine introduce a computational cost that is at most linear in the number of tokens, and runtime is dominated by the quadratic attention component, there is no substantial difference in computational efficiency between the methods.
@@
 LieRE64 0.970G (+0.727%) 5.684G (+1.375%) 20.061G (+1.033%)
-B.8 Validation Losses
+Appendix E Validation Losses
 Table 9: 2D image and 3D video classification Top-1 Validation loss (95% confidence intervals) results.
-All models use 85.1M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Stein et al., 2019) equivalent to DeiT, ∗∗ equivalent to Vivit (spatio-temporal).
+All models use 85.1M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Flanders et al., 2020) equivalent to DeiT, ∗∗ equivalent to Vivit (spatio-temporal).
 Method CIFAR-100 ImageNet-1k UCF101
@@
 LieRE64 1.37 (1.33-1.42) 1.73 (1.70-1.76) 2.64 (2.62-2.67)
-B.9 Basis parameters
+Appendix F Basis parameters scaling
 Table 10: Accuracy Results for Different $\mathrm{LieRE}_{\Theta}$ Parameters, relative to the model size of 85.2M and 88.7M
@@
 UCF101 580608 0.66 % 64 44.68 (44.27-45.09)
-B.10 Validation Accuracy Curves
-Figure 11: CIFAR-100
-B.11 Training Loss Curves
-Figure 12: CIFAR-100
-B.12 Attention Maps
+Appendix G Attention Maps
 We show the attention scores to the CLS token, averaged across heads for every layer.

### v5

@@
 Abstract
-Transformer architectures depend on explicit position encodings to capture token positional information.
-Rotary Position Encoding (RoPE) has emerged as a popular choice in language models due to its efficient encoding of relative position information through key-query rotations.
-However, RoPE faces significant limitations beyond language processing: it is constrained to one-dimensional sequence data and, even with learnable phases, offers limited representational capacity.
-We address these challenges with Lie Relative Encodings (LieRE), which generalizes RoPE to high-dimensional rotation matrices by leveraging their Lie group structure.
-Through extensive evaluation on three image datasets across 2D and 3D classification tasks, LieRE achieves 1.5% improvement over state-of-the-art baselines on 2D tasks and 1% on 3D tasks, while demonstrating superior generalization to higher resolutions.
-Our implementation is computationally efficient, with results reproducible on 4 A100 GPUs in 30 minutes on CIFAR100.
-Our code is available at https://github.com/StanfordMIMI/LieRE.
+Transformer architectures rely on position encodings to model the spatial structure of input data.
+Rotary Position Encoding (RoPE) is a widely used method in language models that encodes relative positions through fixed, block-diagonal, rotation matrices applied to key-query interactions.
+We hypothesize that this inductive bias limits their RoPE’s effectiveness for modalities with high dimensional structure.
+Lie Relative Encodings (LieRE) introduce a principled generalization of RoPE, aimed at increasing the representational capacity of positional encodings in transformers.
+Instead of fixed 2D rotations, LieRE learns dense skew-symmetric matrices (Lie algebra elements), which are then differentiable mapped to form high-dimensional rotation matrices (Lie group elements).
+This results in richer, learnable, and continuous, encodings of both relative and absolute positional information.
+We demonstrate the effectiveness of LieRE on 2D and 3D vision tasks, showing that it generalizes well to higher input resolutions while maintaining computational efficiency.
+The code and checkpoints are publicly available at https://github.com/StanfordMIMI/LieRE.
 Machine Learning, ICML
@@
 This translation-invariant property makes RoPE particularly efficient at capturing position-dependent patterns in text, which has led to its adoption in popular open-source language models such as LLaMA and Mixtral.
-Figure 1: Overview of how LieRE encodes spatial information, where $N$ denotes the number of input dimensions, i.e.
+Figure 1: Overview of how LieRE encodes spatial information, where NN denotes the number of input dimensions, i.e.
 $N=3$ for the 3D image.
-$A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x_{i}\;y_{i}\;z_{i}])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the $i$th patch in the flattened input, where $j$ corresponds to a different patch.
+AA is a learnable skew symmetric matrix and $R_{i}=\exp(A[x_{i}\;y_{i}\;z_{i}])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the $i$th patch in the flattened input, where $j$ corresponds to a different patch.
 $d$ is the head dimension.
@@
 Second row: rotation matrices $\exp(Ax)$ shared across the stem (gray) or different for each layer and head (colorful).
-MHA := Multi-Head-Attention, FF := Feed Forward, L[index] := Layer index, RxH := Rotation Matrix per Head, $\textrm{LieRE}_{\theta}$ := the number of LieRE parameters.
+MHA := Multi-Head-Attention, $F$ := Feed Forward, L[index] := Layer index, RxH := Rotation Matrix per Head, $\textrm{LieRE}_{\theta}$ := the number of LieRE parameters.
 2 Related Work
@@
 5 Experiments
-In order to isolate the effect of changing the position encoding, we use a plain standard transformer backbone and training regime [^1] https://github.com/kentaroy47/vision-transformers-cifar10 with comparable baseline performance on CIFAR100 and ImageNet with (Lee et al., 2022).
-We implement relative position encoding types including LieRE, RoPE-Mixed, Visionllama, and absolute position encodings.
-We use the standard backbone sizes of ViT-Tiny, ViT-B and ViT-L (Dosovitskiy et al., 2020).
-All experiments use RandAugment (Cubuk et al., 2020).
-We avoid using pre-trained weights in order to maximize the comparability of results between methods.
-In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters and use the same default hyperparameters for all experiments (Appendix B.1).
-We evaluate two versions of LieRE, distinguished by the basis matrix tile sizes of 64 and 8, referred to as LieRE64 and LieRE8, respectively.
+We evaluate LieRE across a range of tasks to assess its impact on transformer performance in 2D and 3D vision, spatial reasoning, and resolution generalization.
+Our goal is to isolate the contribution of positional encodings by using consistent architectures, training procedures, and hyperparameters across all methods.
+To this end, we utilize a standard recipe and build on top of the vision transformers repository (Yoshioka, 2024) and verify that our baselines perform similarly to prior work (Lee et al., 2022).
+We avoid using pre-trained weights in order to help reproducibility and comparability of the results between methods.
+We provide confidence interval estimates using bootstrap (B=1000, $\alpha=0.05$).
+We evaluate two versions of LieRE, distinguished by the basis matrix block-diagonal sizes of 64 and 8, referred to as LieRE64 and LieRE8, respectively.
 Notably, a tile size of 2 corresponds to RoPE-Mixed.
-5.1 2D Classification
-For 2D data we evaluate performance on the CIFAR-100 and ImageNet-1k image classification task (Krizhevsky et al., 2009; Deng et al., 2009).
-We further analyse different backbone models sizes and data ablation effects.
-We examine accuracy both CIFAR-100 and ImageNet-1k and compare to absolute (Dosovitskiy et al., 2020) position encoding and recent related work, RoPE-Mixed (Heo et al., 2024), VisionLlama embeddings (Chu et al., 2024).
-For 2D image classification tasks, we demonstrate that the LieRE-based transformer achieves a relative performance improvement in accuracy of 10.0% over standard absolute position encoding (Touvron et al., 2022), 7.3% over RoPE adaptation in VisionLlama (Chu et al., 2024), and 2.2% over RoPE-Mixed (Heo et al., 2024) on CIFAR-100, with similar accuracy trends observed on ImageNet (Deng et al., 2009) (Table 1, Table 2).
-We vary the capacity of the transformer backbone with corresponding to ViT-Tiny, ViT-B and ViT-L, as proposed in (Dosovitskiy et al., 2020).
-Additionally, we analyze the impact of incorporating LieRE on performance across different model sizes on the CIFAR-100 dataset, as shown in Table 6, Figure 8.
-Our results demonstrate that LieRE8 slightly outperforms alternatives across all evaluated model sizes.
+5.1 2D Image Classification
+We begin with CIFAR-100 and ImageNet-1k benchmarks to evaluate LieRE in 2D vision tasks.
+All models use ViT-based architectures trained from scratch with standard data augmentations (RandAugment).
+We compare LieRE to absolute positional encodings, RoPE-Mixed (Heo et al., 2024), and VisionLLaMA (Chu et al., 2024).
+Table 1 shows that LieRE outperforms all baselines.
+On CIFAR-100, LieRE8 achieves a statistically significant relative improvement in top-1 accuracy relatively of 10.0% over absolute encodings, 7.3% over VisionLLaMA, and 2.2% over RoPE-Mixed.
+Similar trends hold on ImageNet (Table  2).
+We also investigate performance across model sizes (ViT-Tiny, Base, Large). As shown in Table 6 and Figure 8, LieRE8 consistently outperforms other methods.
 Figure 4: Performance comparison of different positional embedding methods across varying dataset sizes.
@@
 Both LieRE variants and RoPE-Mixed consistently outperform other methods, with their advantage becoming particularly pronounced in data-scarce scenarios.
-Our experiments reveal a pattern in how learnable relative position encodings perform under limited data conditions.
-Specifically, when we varied the available training data from 20% to 90% of the CIFAR-100 dataset, both LieRE and RoPE-Mixed demonstrated significantly better data efficiency compared to traditional transformer approaches for 2D image classification (figure 4).
-5.2 Synthetic Spatial Classification
+To evaluate robustness in low-data regimes, we perform a data ablation study.
+Figure 4 shows that LieRE variants and RoPE-Mixed maintain significantly higher accuracy than baselines when training on only 20–90% of the CIFAR-100 dataset.
+This highlights LieRE’s data efficiency.
+5.2 Synthetic Spatial Reasoning Task
 Figure 5: The model is asked to identify the direction of the arrow pointed to by the base of the Y. Here said arrow points down. This task requires understanding spatial relationships.
@@
 Figure 5 illustrates an example of this task setup.
+We train the models on 800,000 examples and observe that they generally converge after the first 400,000 examples.
 We verified that none of ChatGPT 4o, Claude Sonnet 3.5 and Gemeni Pro 1.5 are able to solve this task.
-We train the models on 800,000 examples and observe that they generally converge after the first 400,000 examples.
-The only exception to this is absolute position encodings, where all variants were trained on 800,000 examples with no sign of increases in performance.
 Table 3: Comparing different positional encodings on synthetic task on base model (85M) equivalent to DeiT architecture across different resolutions, $108\times 108$, $168\times 168$, $276\times 276$
@@
 We evaluate the models across three different input resolutions ($108\times 108$, $168\times 168$, and $276\times 276$ pixels), revealing differences in scalability and effectiveness.
-While absolute positional encoding shows degraded performance as resolution increases (from 45.1% to 40.1%) and RoPE-Mixed demonstrates strong but degrading performance at higher resolutions (from 100% to 88.6%), both LieRE variants maintain near-perfect accuracy across all resolution scales, with LieRE64 achieving 100% accuracy consistently, suggesting superior resolution invariance of our Lie group-based approach.
+While absolute positional encoding shows degraded performance as resolution increases (from 45.1% to 40.1%) and RoPE-Mixed demonstrates strong but degrading performance at higher resolutions (from 100% to 88.6%), both LieRE variants maintain near-perfect accuracy across all resolution scales, with LieRE64 achieving 100% accuracy consistently.
 Qualitative analysis shows that absolute and visionllama position encoding attend less clearly to the ”Y” token than RoPE-Mixed and LieRE.
@@
 5.3 3D Classification
-In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-Mixed and absolute encoding similar to the previous section (Arnab et al., 2021; Heo et al., 2024).
-For the 3D experiments, we examine video classification performance in the UCF101 dataset (Soomro et al., 2012).
-Again, we did not optimize any hyperparameters for the LieRE model and used the dataloader from (Tong et al., 2022).
+To assess LieRE’s performance on 3D data, we use the UCF101 video classification benchmark (Soomro et al., 2012).
+All models use a ViT-style backbone with 3D patch tokenization, trained from scratch with no hyperparameter tuning and the dataloader from (Tong et al., 2022).
 The full set of hyperparameters may be found in appendix B.1.
 We observe a relative accuracy improvement of the LieRE-based transformer of up to 15.1% compared to absolute position embeddings and at least 1.5% compared to RoPE-inspired position encodings (table 1).
-5.4 LieRE Capacity
-LieRE adds a small amount of capacity to the model ($580k$ parameters 0.68% for the ViT-B backbone we use for most experiments), leading to the natural question of how helpful the marginal capacity is.
-To answer this question, we vary the density of the learned basis and the learnability between heads and layers.
-We vary capacity by enforcing a block-diagonal structure on the skew-symmetric basis.
-Varying the block size allows us to approach a dense matrix (LieRE64 for our models).
-Recall that using $2\times 2$ blocks recovers exact commutativity and is identical to RoPE-Mixed.
-As RoPE-Mixed is sensitive to the initialization of the weights, we initialize all generator weights mirroring the original presentation of RoPE-Mixed (Heo et al., 2024).
-We found that RoPE-Mixed is more sensitive to the initialization method than LieRE as seen in appendix table 7.
-In figure 6 we evaluate accuracy versus block dimension (Appendix Table 10).
+5.4 LieRE Capacity and Parameter Efficiency
+LieRE introduces minimal overhead–—only 580k additional parameters (0.68% for ViT-B)–—yet offers a flexible mechanism for increasing representational capacity.
+To explore the impact of this marginal capacity, we vary the density of the skew-symmetric basis and examine parameter sharing across heads and layers.
+We control capacity via imposing a block-diagonal structure on the basis matrices.
+Smaller blocks (e.g., $2\times 2$) replicate RoPE-Mixed, while larger blocks increase expressivity, with LieRE64 using a fully dense basis.
+We adopt RoPE-Mixed’s initialization for fair comparison.
+We observe that RoPE-Mixed is more sensitive to initialization than LieRE (Appendix Table 7).
+Figure 6 shows performance as a function of block size.
+Both in 2D (CIFAR-100) and 3D (UCF101), accuracy improves with block size, peaking around $8\times 8$—–suggesting this is a sweet spot between capacity and regularization.
 Figure 6: Performance varies with skew-symmetric basis learnable dimensions, referred to as LieRE parameters.
@@
 5.613G 68.3 69.4 69.5
-Table 4 presents the impact of sharing LieRE parameters across attention heads and layers on CIFAR-100 classification performance for RoPE-Mixed, LieRE64 and LieRE8 .
-We evaluate whether learning separate positional encodings for each attention head and layer provides performance benefits.
-We observe that learning across heads and layers yields superior performance followed by learning across layers.
-5.5 Patch shuffling
+We also assess the effect of parameter sharing on CIFAR100 (Table 4).
+Learning LieRE parameters independently per head and per layer yields the best results.
+Sharing across layers or heads reduces accuracy, but still outperforms RoPE-Mixed.
+5.5 Patch shuffling: Measuring Positional Dependency
 Shuffling patches and frames allows us to see how much the model is able to use the positional information in its inputs.
 A model whose architecture does not allow/encourage the use of positional information will converge to a representation similar in spirit to a bag-of-words, where the relative locations of pixels/voxels do not matter.
-A greater dropoff in accuracy during shuffling is indicative that the model more heavily utilizes positional information.
+A greater drop-off in accuracy during shuffling is indicative that the model more heavily utilizes positional information.
 We evaluate models using the decline in accuracy when evaluating on shuffled patches.
-We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE models are more capable at using positional information.
+We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE models rely more on positional information as expected.
 The complete results are displayed for CIFAR-100 and table for UCF101 (table 5).
@@
 Method Before After Drop(%) Before After Drop(%)
-[…]
+$\uparrow Shuffling\downarrow$ $\uparrow$ $\uparrow Shuffling\downarrow$ $\uparrow$
 Abs. Pos. E. 63.9 19.6 69.3 40.9 39.5 0.0
@@
 5.6 Multi-resolution Classification
-In this section we compare the ability of methods to generalize to image resolutions not seen during training.
-We evaluate two training recipes inspired by (Heo et al., 2024).
+In this section we compare the ability of methods to generalize to image resolutions not seen during training. We evaluate two training recipes inspired by (Heo et al., 2024).
 The first recipe matches the rest of the paper and consists of training the models on images of size $224\times 224$ for 200 epochs.
@@
 6 Conclusion
-In this paper, we introduced Lie group Relative position Encodings (LieRE), a position encoding that can effectively encode relative position information for attention mechanisms across modalities and input dimensionalities.
-Through experiments on 2D image classification (CIFAR-100, ImageNet-1k) and 3D video classification (UCF101), we demonstrated that LieRE achieves better performance compared to existing positional encoding methods.
-Beyond improving accuracy, LieRE also exhibits data and compute efficiency.
-On CIFAR-100, LieRE requires 3.5 times less training computed to match the performance of the baseline model with absolute position encodings.
-Furthermore, LieRE can outperform the baseline trained on the full dataset while using only 70% of the training data, highlighting its data efficiency.
-The key advantages of LieRE include its simplicity, flexibility, and ease of adaptation to new modalities.
-By requiring no changes to the tokenizer other than outputting positions and no other code changes, LieRE may provide a unified and efficient approach for transformers to process and learn from various data modalities within a single architecture.
+We proposed Lie group Relative position Encodings (LieRE), a positional encoding method that modifies the attention method via dense, learned, high-dimensional rotation matrices.
+As compared to the more widely used block-2D rotation matrices typically used to encode positions, dense rotation matrices can encode both relative and absolute positional information and allow a greater portion of the model’s learnable capacity be allocated to spatial reasoning.
+Experiments on 2D image classification (CIFAR-100, ImageNet-1k) and 3D video classification (UCF101) show that LieRE consistently outperforms existing positional encoding methods.
+Our analysis indicates that LieRE-based ViTs effectively leverage spatial reasoning capabilities unavailable to transformers using only absolute positional encodings.
+In addition to accuracy gains, LieRE offers notable data and compute efficiency.
+Its simplicity, flexibility, and strong capacity to learn spatial structure make it broadly applicable.
+With no tokenizer modifications beyond position output and no additional architectural changes, LieRE provides simple approach for controlling the amount of positional information into transformers.
 7 Limitations
-While LieRE shows promising results across multiple modalities and input dimensionalities, there are a few limitations worth noting.
-Our method is specifically designed to modify the inner product, making it compatible with most attention schemes, including original attention and linear attention.
-However, this specificity may limit its applicability to other architectures, like convolutional neural networks, that do not rely on inner product-based attention mechanisms.
-Future work could explore adaptations of our method to a wider range of architectures.
-Furthermore, the current formulation of our method is designed to encode vector positions in $\mathbb{R}^{d}$.
-While this is sufficient for many applications, it may not be directly applicable to tasks that require encoding poses in $SE(3)$ like robotics.
-Further research may be necessary to adapt our method to effectively handle such representational requirements.
-Despite these limitations, we believe that our work provides much-needed insight into how to improve model performance and reduce training costs by encoding relative position information across various modalities and input dimensionalities.
+While LieRE shows promising results for 2D and 3D inputs, several limitations are worth noting.
+For 1D input, LieRE reduces to RoPE with learnable phases (proof in appendix A).
+Our method is designed to modify the inner product, making it compatible with most attention mechanisms, including standard softmax attention and linear attention.
+However, this may limit its applicability to other architectures—–such as convolutional neural networks—–that do not rely on the attention mechanism.
+Future work could adapt the method to a broader range of architectures.
+The current formulation encodes vector positions in $\mathbb{R}^{d}$.
+While sufficient for many applications, it may not directly apply to tasks that require pose encoding in $SE(3)$ (e.g., robotics).
+Lastly, in its current implementation, LieRE relies on the accuracy and numerical stability of the matrix exponential in PyTorch.
+Future work may explore more efficient and robust implementations or approximations of this operation.
+Despite these limitations, we believe our approach provides valuable insight into improving model performance and reducing training costs by encoding relative positional information across various input dimensionalities.
 Impact statement
@@
 We would like to thank Aradhana Sinha for suggesting the shuffling experiment and providing early feedback on the paper.
-Sophie Ostmeier was supported by the German Research Foundation (DFG), Walter-Benjamin fellowship (ID: 517316550).
+Furthermore, lucidrains suggested adding the proof of that in the 1D setting, LieRE has identical representational capacity to RoPE up to the learnable basis.Sophie Ostmeier was supported by the German Research Foundation (DFG), Walter-Benjamin fellowship (ID: 517316550).
 This project was supported by Google Cloud and Azure credits.
+This work was also supported in part by the Medical Imaging and Data Resource Center (MIDRC), which is funded by the National Institute of Biomedical Imaging and Bioengineering (NIBIB) under contract 75N92020C00021 and through the Advanced Research Projects Agency for Health (ARPA-H).
 Appendix A Equivalence of RoPE and LieRE in one Dimension
@@
 The ImageNet experiments were trained on 8xL40 GPUs and all took less than 2 days and 5 hours of runtime including time lost due to preemption and resource sharing.
-We use a cosine learning rate schedule with an initial learning rate of $41E-$ 4 and train for 200 epochs.
+We use a cosine learning rate schedule with an initial learning rate of 1E41E-4 and train for 200 epochs.
 We use an effective batch size of $512$.
@@
 B.3 3D Video Classifications
-The 3D classification experiments were conducted on either $1008\times A$ 100 40GB GPUs or $1004\times A$ 100 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size.
-Similar to 2D classification, we use an initial learning rate of $41E-$ 4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $2\times 16\times 16$ on the randomly cropped and resized $8\times 224\times 224$ video/image.
+The 3D classification experiments were conducted on either $8\times A100$ 40GB GPUs or $4\times A100$ 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size.
+Similar to 2D classification, we use an initial learning rate of 1E41E-4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $2\times 16\times 16$ on the randomly cropped and resized $8\times 224\times 224$ video/image.
 We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$.
@@
 In order to understand the sensitivity of the methods to the initialization, we explore several scaling factor for the weights initialization. The results are presented in table 7.
-Table 7: Paired Z-test between $2\pi$ and $1$ initialization for LieRE8 and RoPE-Mixed.
+Table 7: Paired Z-test between $2\pi$ and 11 initialization for LieRE8 and RoPE-Mixed.
 Metric LieRE8 ($2\pi$ vs 1 init) RoPE-Mixed ($2\pi$ vs 1 init)
@@
 Appendix C Python implementation of LieRE rotation matrix computation
-basis_raw_params = nn.Parameter( torch.rand( input_dimensionality, head_dim, head_dim, ) * 2 * math.pi # optional, inspired from RoPE-Mixed paper ) upper_triangle = ( torch.triu(basis_raw_params, diagonal=1) ) skew_bases = upper_triangle - torch.transpose(upper_triangle, -1, -2) in_basis_positions = ( positions.reshape(list(positions.shape) + [1] * 2) * skew_bases ) rotation_log = torch.sum(in_basis_positions, dim=-3) rotation = torch.matrix_exp(rotation_log.to(dtype=torch.float32)) rotation = rotation.to(dtype=positions.dtype)
+basis_raw_params = $n$.Parameter( torch.rand( input_dimensionality, head_dim, head_dim, ) * 2 * math.pi # optional, inspired from RoPE-Mixed paper ) upper_triangle = ( torch.triu(basis_raw_params, diagonal=1) ) skew_bases = upper_triangle - torch.transpose(upper_triangle, -1, -2) in_basis_positions = ( positions.reshape(list(positions.shape) + [1] * 2) * skew_bases ) rotation_log = torch.sum(in_basis_positions, dim=-3) rotation = torch.matrix_exp(rotation_log.to(dtype=torch.float32)) rotation = rotation.to(dtype=positions.dtype)
 Appendix D Compute efficiency
-We demonstrate that LieRE-based transformer achieves comparable performance to the Absolute Position Embedding baseline (DeiT III (Touvron et al., 2022)) on CIFAR-100 with fewer training epochs.
-This represents a notable advancement over recent methods such as VisionLlama and RoPE-Mixed (Chu et al., 2024; Heo et al., 2024).
-Figure 9 illustrates that LieRE enables a 3.9$\times$ reduction in training compute while maintaining the accuracy achieved by absolute position encodings after 200 epochs.
+[…]
 Figure 9: The LieRE spatial encoding allows the model to match the performance of absolute position encodings with substantially less training time.